A Complete Bibliography of Publications in IEEE Transactions on Software Engineering (1970–1979)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

05 April 2023 Version 1.01

Title word cross-reference

3 [SB77]. → [MW79].

1 [BRGP78].

2 [Ano76a].

3 [Ano77a]. 360 [HP78].

4 [Ano78a].

5 [Ano79a].

Abstraction [ZL78]. Abstractions [GY76a, LZ75]. Access [Cer77, Gli79, Har77a, HKN79, JL76, MA79]. Achieve [Ano76m]. Acknowledgment [Ano76b, Ano77b, Ano78b, Ano79b]. Address [Smi77]. Advance [Ano76c, Ano78c, Ano78d, Ano78e, Ano79c, Ano79d, Ano79e]. advertisement [Ano76d, Ano77d, Ano77e, Ano77f]. Aid [Ham77, Rei77, Sto77a]. Aided [BBD77, TH77]. aiding [RKC75]. aids [BMU75]. Aladdin [Fai79]. Algebra [Sto77b, Pai75]. Algorithm [Mis78c]. Algorithms [Rob79, Smi78, Sto77a, Zwe77]. Allocation [Boc78, SB78]. Allow [SKB77]. Alphanumeric [SF77]. Alphard [WLS76]. Alternatives [Ano77c]. Analysis [AC79, Ano78f, Ano78v, CHT79, Chi78, CC77, Els76, Fra78, Goe80, GM79, Har77b, How76, KSY77, Kod78a, LCC76, Rob79, RS77, Ros77, SW78, SH79a, SH79b, Smi77, TH77, End75, SV75]. Analytic [CBDU75]. Analytical [Mar79]. Analyzing [Wat79].

Announcement [Ano78d, Ano79c, Ano79d, Ano79e]. Announces [Ano78-29, Ano78-30]. Announcing [Ano76d, Ano77c, Ano77d, Ano77e, Ano77f, Ano78f, Ano78v, Ano79f, Ano79g, Ano79h]. Anomaly [Hua79]. Application [And79, CC78, CK78, Fra78, LCC76, Mus75]. Applications [Ben79, GY76b, NH79, OB79, Rou79]. Applied [Mcc75a]. Apprentice [RS78, HS75]. Approach [BBD77, BJ76, GS78, LLS81, Mis78a, Wit75]. Approximate [Mar79]. architecture [Wul75]. Argument [Rom77]. Artificial [Kli77]. Aspects [BSW79, Mis78b, RDB75]. Assemblers [Spi76]. Assembly [Fai79]. Assertion [Fai79]. Assertions [Wor79, GW75]. assessment [Wir75]. Assignment [Gri78]. ASSIST [HP78]. ASSIST-V [HP78]. Assistant [Mor79]. Associations [Gho76]. Associative [Smi78]. Asynchronous [ABSS78, Rob79]. Attribute [KSY77]. Auditing [Ano76y, LSM78]. Authentication [GL79]. Author [Ano75a, Bak76]. authors [Ano75k, Ano77-39, Ano78-39, Ano79-39]. Automated [IB77, RHC76, BMU75, RH75]. Automatic [DM77, LSM78, MS76, Pol79, Sto77b]. Automating [RT78]. Average [Cer77]. Avoidance [RL79].

Back [Ano75b, Ano75c, Ano75d, Ano75e]. Backtracking [GY76a]. Base [Gho76, GT76, KSY77, Rou79, CBDU75]. Based [SF77, ZL78]. Bases [DDL78, MM79]. Basis [RY76]. Batch [Ell78]. Be [DDL78]. Behavior [RWS^{+}78]. Better [Wal77]. Between [BJ76, Old77, SB78]. Binary [Ben79]. Blocking [CP79]. Books [Ano78g]. bottom [Mcc75b]. bottom-up [Mcc75b]. Building [WG76].

Cache [Smi78]. Call [Ano76i, Ano77m, Ano77n, Ano77o, Ano77p, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ryd79]. Can [DDL78]. Capabilities [EB79, Gli79]. Capability [KS78, ZL78]. Capability-Based [ZL78]. Capacities [CHS77]. Capacity [CP79]. Case [Ano76m, BRGP78, CM79, CS79, LP79, Mcc75a]. Cassette [Cer77]. causes [End75]. Challenge [DN78]. chapter [Ano76z]. Character [RH76]. Checking [Sto77b]. Chi [LLS75]. Chi/OS [LLS75]. Clarity [Gor79a, Gor79b]. Closed [GM79, TW79]. CLU [LS79]. Clusters [Kat79]. Code [Har79, Roc75]. Collection [Ano78u, SH79a, SH79b]. COLUMBUS [Wit75]. Comments [LLS81, Tom76]. Commercial [Els76]. Common [LLD79]. Communicating [Ros77]. Communication [CHS77, DJ77, Sil79]. communications [SS75]. Comparative [Smi78]. Comparison [Nut77]. Compcon77 [Ano77q]. Competing [Goe80, SW78]. Compile [AC79, Boc78]. Compile-Time [AC79]. Compiler [BE77, Ham77, Har77b, Har79, Sam77, Tan76]. Compiler-Compiler [Tan76]. complete [Ano76v, Ano76w, Ano76x, Ano77-34, Ano77-35, Ano77-36, Ano77-37, Ano77-38, Ano79-36]. Complexity [Che78, CSM^{+}79, McC76, Woo79, WHH79]. Component [LCC76]. Compromised [DDL78]. computation [SB75]. Computations [BK76, Mis78b, Pra78]. Computer [Ano76d, Ano76r, Ano76n, Ano76s, Ano76o, Ano76t, Ano76p, Ano76u, Ano76q, Ano76y, Ano76z, Ano77c, Ano77d, Ano77e, Ano77f, Ano77-28, Ano77v, Ano77-29, Ano77w, Ano77-30, Ano77x, Ano77-31, Ano77y, Ano77-32, Ano77z, Ano77-33, Ano77-27, Ano78g, Ano78-31, Ano78w, Ano78-32, Ano78x, Ano78-33, Ano78y, Ano78-29, Ano78-34, Ano78z, Ano78-35, Ano78-27, Ano78-36, Ano78-30, Ano78-28, Ano79f, Ano79g, Ano79h, Ano79-29, Ano79w, Ano79-30, Ano79x, Ano79-31, Ano79y, Ano79-32, Ano79z, Ano79-33, Ano79-27, Ano79-34, Ano79-28, BBD77, CHS77, Gho76, HW77, MW78a, Sto77b, TH77, Ano76v, Ano76w, Ano76x, Ano77-34, Ano77-35, Ano77-36, Ano77-37, Ano77-38, Ano79-36, MBGK75]. Computer-Aided [BBD77, TH77]. Computers [Ano76j]. Computing [Sto77b]. Concepts [BZ79]. Conceptual [Rou79]. Concurrency [BRGP78, BSW79, Sto79]. Concurrent [Han77, KY75, Sch79, Bri75, SKB77]. Conditional [EB79]. Conditions [ABSS78, Van79]. Conference [Ano78u]. configuration [WD75]. Consistency [Sto79, Sto77b]. Constrained [Kun78]. Constrained-Path [Kun78]. Construct [BBS^{+}79, For78]. Constructing [BK76, Ryd79]. Construction [Mis78c, WLS76]. Constructive [Weg77]. contents [Ano75p, Ano75q, Ano75r, Ano75s, Ano76e, Ano76f, Ano76g, Ano76h, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n]. Context [HKN79, SO77]. Contraction [Par79]. Contrast [Old77]. Control [Ada77, BRGP78, BSW79, GK78, GY76a, HKN79, MA79, Pra78, Sto79, WHH79, Roc75]. Controlling [CS78, JL76]. Convention [Tow79]. Conversion [Sto77b]. Copies [Sto79]. copyright [Ano79-35]. Corporate [Coo78]. Correct [Gri76, Gri77]. Correction [Ano76k, GG75a, Gri77, SH79a, Tai78]. Correctness [Bab79, Gri76, Gri77, Lam77]. Correspondences [AC79]. Corrigendum [Elg76a]. Cover [NH79, Ano75l, Ano75m, Ano75n, Ano75o]. Critical [For78, LLS81, Sto78a]. CSDL [Rou79]. Current [Gri76, Gri77]. Cycle [CS78, DN78, MW78b]. Cyclic [Fra78].

Data [AC79, Ano77t, BBS^{+}79, Chi78, Cla76, DDL78, Far77, Gho76, GT76, Hua79, IB77, JL76, KSY77, Kod78a, LP79, MM79, MS76, RHC76, Rip77, Rou79, SH79a, SH79b, Smi77, Sto79, Tow79, ZL78, CBDU75, GM75, GG75a, GG75b, LZ75]. Data-Base [KSY77]. Database [Ben79, BSW79, CC78, CK78, DLSW78, HY79, HKN79]. Databases [BRGP78]. Deadlock [MM79, RL79]. Debugging [BBS^{+}79, Fai79, Ott79]. Decision [DJ77, TW79]. Declaration [Tow79]. Defense [DN78]. Defining [HZ76, Led77]. Definition [MS77, RS77, Rou79]. Definitions [Mis78a]. Deletions [Knu77]. Department [DN78]. Dependent [HKN79]. Derivation [Gri76, Gri77, LP79]. Descent [CSA79]. Description [Sam77, SB77]. Descriptions [Sha78]. Design [And79, BE77, BP78, CM79, Cho78, CH79, GS78, GT76, LLS81, Par76, PHW76, Pra78, RWS^{+}78, Rou79, Sch79, BMU75, GH75, Voi75]. Designer [Mor79]. Designer/Verifier [Mor79]. Designing [Par79]. Detection [Hua79, MM79, RL79, KY75]. Developer [MW78b]. Development [Ano76m, BJ79, Dal77, DV77, HP78, Mil76, Par76, Rei77, BT75]. Dialogues [Kat79]. Differential [BRH79]. Digital [Ano78-29, Ano78-30, Cer77]. Digraphs [NH79]. Direct [CC76]. Direct-Execution [CC76]. Directed [Weg76]. Discrete [Kod78b]. Discussion [Goe80]. DISSECT [How77, How78a]. Distributed [BRGP78, CM79, Chi78, Gli79, HY79, MM79, Sil79, Sto78a, Sto79]. Distributing [Gho76]. Distribution [ZH79]. Do [Ano76m]. Documentation [TH77]. doesn’t [Ano76v, Ano76w, Ano76x, Ano77-34, Ano77-35, Ano77-36, Ano77-37, Ano77-38, Ano79-36]. Down [CH79, Tow79, Mcc75b]. Drawings [WS79].

Dreams [MW79]. Driven [Fai79]. Dual [Bok79]. Due [CP79]. During [RWS^{+}78]. Dynamic [Ano76k, BS76, Bok79, Cla79, GR78, GIL78, OB79, SKB77].

Ease [Par79]. Editor [Ano75f, Ano75g, Ano75h, Ano77r, Ano77s, Ano78s, Ano79u, Ano79v, Ano79t]. Editorial [Ano76l, Ano76m, Ano77t, Ano77u, Ano78t, Ano78u, Hal79, Mer78, Ano75j, Ste75]. Efficient [BS79, Smi77]. Empirical [How78b, Put78]. Encoded [SF77]. Engineering [Alf77, BBD77, HW77, IB77, Wil76, Ano76v, Ano76w, Ano76x, Ano77-34, Ano77-35, Ano77-36, Ano77-37, Ano77-38, Ano79-36, Chu75, Cul75, Mcc75a, Ano78u, Ano76a, Ano76e, Ano76f, Ano76g, Ano76h, Ano77a, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78a, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n]. enhancement [BT75]. Environment [BE77, Chi78, HP78, Tom76, Bak75]. Environments [RH76, Tow79]. Equations [BRH79]. Equivalence [ABSS78]. Erratum [Elg76a]. Error [SH79a, SH79b, Tai78]. errors [End75]. ESP [SB77]. Estimates [Ott79]. Estimating [Put78]. Euclid [Wor79]. Evaluating [BZ79, CSA79]. Evaluation [And79, BE77, BRH79, Cer77, CHT79, How77, How78a, RH75]. Even [DDL78]. Everything [Ano76m]. Evolution [DM77, RW76]. Example [BK76]. Exception [LS79]. Exclusion [HS78]. Execute [Cla76]. Execution [CC76]. Exercise [Pol79]. Experience [Han77, BMU75, LLS75]. Experiment [CH79, GT76, SH79a, SH79b, Woo79]. Experimental [GIL78, Fai75]. Experimentation [Rei77]. Explicit [Tow79]. Expressions [Sha78]. Extendable [BBD77]. Extending [SKB77]. Extension [JL76, Par79]. extraction [Weg75].

Facility [Sam77, Fai75]. Factors [Sto78a]. Fall [Ano77q]. Fallibility [GY76b]. Families [Par76]. Family [PHW76]. Fault [GS78, LLS81, KY75, Ran75]. Favor [Rom77]. File [LLD79]. Find [Mer78]. Finite [CP79, Cho78]. Finite-State [Cho78]. Fixed [RH76]. Floating [MS76]. Floating-Point [MS76]. Flow [Hua79, Sha78, Sto77a, WHH79]. Flowcharts [Kod78b]. Flowgraphs [Kod78a]. Form [HW77, Ano79-35]. Formal [BSW79, CH78, Mis78a, Noo75]. Format [AC79]. Forms [GR78]. Fortran [WD75, RW76, RH76, WD75]. founded [Weg75]. Free [Chi78]. Frequency [ZH79]. Front [Ano75i, Ano75l, Ano75m, Ano75n, Ano75o]. Fully [BRGP78]. Fuzzy [CK78].

Game [HW77]. Gannon [LLS81]. Gateway [BFL^{+}76]. General [Ada77, Har79, Mar79, Put78, Tan76]. General-Purpose [Har79, Tan76]. Generalized [For78]. Generate [Cla76, Ell78]. Generation [GMO76, Har79, MS76, RHC76]. Generator [CL77, Gai77]. GO [Elg76a, Elg76b]. Goal [Weg76]. Goal-Directed [Weg76]. Grammar [GR78]. Graph [Ryd79]. Graphs [Pai77, Pai75]. group [SS75]. Guest [Ano75j, Ano76l, Ano76m, Ano77t, Ano77u, Ano78t, Ano78u, Hal79, Mer78, Ste75].

H [Ano79-37]. Halstead

[Ano79-37, CSM^{+}79]. Handling [LS79]. hardware [Wul75]. hardware/software [Wul75]. High [CC76, Wor76]. High-Level [CC76]. High-Resolution [Wor76]. Higher [HZ76]. history [Mcc75a]. Hut [HW77].

IBM [HP78]. Ideas [Gri76, Gri77, Ros77].

IEEE [Ano76d, Ano76r, Ano76n, Ano76s, Ano76o, Ano76t, Ano76p, Ano76u, Ano76q, Ano76z, Ano77d, Ano77e, Ano77f, Ano77-28, Ano77v, Ano77-29, Ano77w, Ano77-30, Ano77x, Ano77-31, Ano77y, Ano77-32, Ano77z, Ano77-33, Ano77-27, Ano78f, Ano78g, Ano78v, Ano78-31, Ano78w, Ano78-32, Ano78x, Ano78-33, Ano78y, Ano78-29, Ano78-34, Ano78z, Ano78-35, Ano78-27, Ano78-36, Ano78-30, Ano78-28, Ano79f, Ano79g, Ano79h, Ano79-35, Ano79-29, Ano79w, Ano79-30, Ano79x, Ano79-31, Ano79y, Ano79-32, Ano79z, Ano79-33, Ano79-27, Ano79-34, Ano79-28, Ano76a, Ano76e, Ano76f, Ano76g, Ano76h, Ano77a, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78f, Ano78v, Ano78a, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79i, Ano79j, Ano79k, Ano79l]. IEEE [Ano79m, Ano79n]. If [Ano76v, Ano76w, Ano76x, Ano77-34, Ano77-35, Ano77-36, Ano77-37, Ano77-38, Ano79-36]. Illustration [Gri76, Gri77]. Image [Ano78-29, Ano78-30, Har77a]. Implementation [BJ79, DJ77, For78, HS78, RGH78, RT78, ZL78]. Implementations [Kat79]. Implementing [CL77]. implication [Pai75]. Improvement [Ano76k, BS76, Mas79]. Improvements [Gor79a]. Improving [CSA79]. Impure [Sto77b]. Increase [Woo79]. independent [WD75]. Index [Ano76a, Ano77a, Ano78a, Ano79a, Ano75a]. Induction [RY76]. inductive [GW75]. Industrial [PB78]. Influence [Els77]. Informality [BGW78]. Information [Ano75k, Ano77-39, Ano78-39, Ano78-37, Ano78-38, Ano79-39, Ano79-38, Boo79, SF77, TH77]. Ingres [Sto79]. Initial [RS78]. Injection [GK78]. insertion [PS75]. Instrumentation [Hua79]. Integer [RH76]. Intelligence [Ano78f, Ano78v]. Intelligent [BFL^{+}76]. Interactive [CC76, DHI78, Ell78, Far77, GLB75]. interfering [Mck75]. International [Ano78u]. Interpreter [Fai79]. Introducing [Moo75]. Introduction [BP78, WLS76]. Invariants [Kat78]. iteration [Moo75]. Iterative [Ada77, BT75].

Justification [Gor79b].

Keys [SF77]. Kongsberg [PB78].

Language [CC76, Fai79, GH75, JL76, Old77, Rei77, Ros78, Ros77, Rou79, Sch79, SB77, Bri75, Wir75]. Languages [Boo79, Tai78, Mck75]. Large [DK76, BMU75, RH75]. large-scale [BMU75]. Learn [Ano76m]. Legality [Wor79]. Level [CC76, Coo78, Old77]. library [Ano76v, Ano76w, Ano76x, Ano77-34, Ano77-35, Ano77-36, Ano77-37, Ano77-38, Ano79-36]. Lie [DDL78]. Life [CS78, DN78, MW78b]. Limits [Ano77c]. Lisp [Cla79, Moo75, RS78]. List [AC79, Cla79]. List-Format [AC79]. Load [Sto78a]. Locality [Ano76k, BS76, Mas79]. Locking [MM79]. Logic [MW78a]. Logical [CC78, Gho76, RL79]. Loop [Mis78b, Pra78, Wat79, BM75].

Machine

[Ano78f, Ano78v, DJ77, Sam77, WD75]. Machines [Cho78, CH78]. Macro [Put78, Tan76]. Macroprocessor [Ell78]. Main [Smi78]. Maintenance [CSM^{+}79]. Man [Tan76]. Management [Ano76y, CS78, Chi78, Coo78, Dal77, DN78, IB77, MW78b, Mer78, RGH78, SKB77]. Managers [KS78]. Mapping [Smi78]. matter [Ano75b, Ano75c, Ano75d, Ano75e, Ano75i]. Maurice [Ano79-37]. McCabe [CSM^{+}79]. Measure [Gor79b, McC76, WHH79]. Measurement [Mas79, Rip77].

Measurements [Cla79]. Measures [Old77]. Measuring [CSM^{+}79, Gor79a]. Mechanism [BRGP78, MS77]. Mechanized [Rom77]. Memory [Ano76k, Ano79-37, BS76, BS79, Boc78, Cer77, GK78, Mas79, Smi77, Smi78]. Message [And79]. Method [Fra78, Mar79, Wat79]. Methodologies [GY76b]. Methodology [Alf77, BP78, HZ76, PB78, Chu75]. Methods [Mas79, Smi77, Weg77]. Metrics [CSM^{+}79]. Microprocessor [CC76]. Migration [GL79]. millions [Ano76j]. Minicomputer [CL77]. Minimal [PHW76]. Model [ALS78, BE77, GT76, TW79, SS75]. Modeled [Boo79, Cho78]. Modeling [CH78, RWS^{+}78, SB75]. Models [Goe80, SW78, CBDU75]. Modern [GY76b]. Modification [DM77]. Modula [And79]. Modular [Han77]. Modularity [BZ79]. monitoring [MBGK75]. monitors [RKC75]. Multiattribute [LLD79]. Multidimensional [Ben79]. multidisciplinary [Voi75]. Multikey [LCC76]. Multilevel [ALS78]. Multiple [Gri78, KSY77, Sto79]. Multiple-Attribute-Tree [KSY77]. Multiprocess [Lam77]. Multiprocessor [Han78, Rob79, Sto77a]. Multiprogramming [GK78]. Must [Mer78]. Mutual [HS78].

Need [Ano76m]. Network [Gho76, Sto77a, MBGK75, Han78]. Networks [GM79, Mar79, Spi79, TW79]. Nonlinear [Gai77]. Notation [Led77]. Note [Ano79t, Kun78]. Notice [Ano77r, Ano77s, Ano78s, Ano79u, Ano79v, Ano75f, Ano75g, Ano75h]. Number [CL77, Gai77]. Numbers [Sto77b].

Object [GL79]. Observations [GY76b]. Operating [Chi78, Ell78, GIL78, Nut77, PHW76].

Operation [BS79]. Operations [Ano76y]. Operators [ZH79]. Optimal [GS78, LLS81, RKC75]. Optimization [Boo79, Kat78]. Optimizing [Har79]. Order [HZ76]. Organization [KSY77]. Organize [Ano76z]. Overview [BFL^{+}76]. own [Ano76z].

Paper [LLS81]. Papers [Ano76i, Ano77m, Ano77n, Ano77o, Ano77p, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano78-29, Ano78-30, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s]. Parallel [Bab79, BE77, Boc78, Gho76, MA79, Nut77]. Parsers [CSA79]. Partial [BRH79]. Partitioning [Pai77]. Pascal [Bri75, SKB77, Wir75]. Path [How76, Kun78, NH79]. Paths [GMO76]. Pattern [Ano77t, Ano78f, Ano78v, Far77, SB77]. Patterns [Kli77]. Perceived [DHI78]. Perfection [Ano76m]. Performance [BRH79, Boo79, GT76, RGH78, SB75]. Periodic [GM79]. Perspectives [Rip77]. PFortran [WD75]. Physical [Zwe77]. Picture [WG76]. Picture-Building [WG76]. PL [Els76, Els77, MS77, ZH79]. PL/I [Els76, Els77, MS77, ZH79]. placement [RKC75]. Plan [Ano76d, Ano77d, Ano77e, Ano77f]. Point [MS76]. Policies [CC77]. Polish [CS79]. Poor [Tan76]. Portable [RH76, WD75]. practical [BT75]. Predicting [SS75]. Presence [HKN79]. Preserve [Knu77]. preventing [Mck75]. Principal [LCC76]. Principles [Mis78a]. priority [PS75]. Privileges [Gli79]. Probabilistic [Boo79]. Problem [BJ76, Kun78, Put78, SO77, Woo79, GM75]. Problem-Reduction [BJ76]. Problem-Solving [SO77]. Problems [GMO76, NH79]. Process [MW78b]. Processes [Boc78, Sch79, SB78]. Processing [Alf77, Ano78-29, Ano78-30, BE77, BBS^{+}79, Boo79, Har77a, HY79, RH76, TH77]. Processor [Bok79, Nut77, Sto78a, Tan76]. Production [Led77, Rom77, Tom76, Bak75]. Productivity [Che78, SS75]. Professionals [Ano78g]. Profiles [Els77]. Prog [Fra78]. Program [Ano76c, Ano78c, Ano78d, Ano78e, Ano79d, Ano79e, BGW78, BBS^{+}79, Che78, CS79, DM77, Els77, GMO76, Gor79a, Gor79b, Han78, HW77, How78a, How78b, Hua79, Kat78, Kun78, LP79, Mas79, NH79, Pai75, Pai77, Par76, Pol79, RHC76, RY76, Rip77, Ryd79, Voi75, Wal77, Weg76, Weg77, WHH79, Fai75, GLB75]. program-testing [Fai75]. Programmer [Che78, RS78]. Programming [DK76, Elg76a, Elg76b, Els77, GY76a, GY76b, Han77, MW78a, Mis78a, Sch79, Tai78, Tom76, Van79, Wil76, Bak75, Bri75, GH75, HS75, Mcc75b, Mck75, Noo75, Pai75, SS75, SV75, Wir75]. Programming-in-the-Large [DK76]. Programming-in-the-Small [DK76]. Programs [Bab79, BS79, BK76, BJ76, CM79, CHT79, Cla76, DM77, Els76, Gri76, Gri77, Ham77, Lam77, MW79, MA79, Wat79, WLS76, ZH79, BM75, BY75, End75]. Project [CS78, HW77]. Proof [HA79]. Proofs [Gri76, Gri77, Mis78a]. Properties [LLD79]. Property [Weg75]. Protection [HKN79]. Protocol [Har77a]. prover [Moo75]. Proving [Bab79, BM75, BJ76, Lam77]. PSA [TH77]. Pseudorandom [CL77, Gai77]. PSL [TH77]. PSL/PSA [TH77]. Psychological [CSM^{+}79, Rei77]. Publications [Ano76r, Ano76s, Ano76t, Ano76u, Ano77-28, Ano77-29, Ano77-30, Ano77-31, Ano77-32, Ano77-33, Ano78g, Ano78-31, Ano78-32, Ano78-33, Ano78-34, Ano78-35, Ano78-36, Ano79-29, Ano79-30, Ano79-31, Ano79-32, Ano79-33, Ano79-34]. Pure [Moo75]. Purpose [Har79, Tan76].

Qualitative [Gor79b]. Quality [DHI78].

Quantitative [Ott79, RDB75]. Quarterly [Ano78f, Ano78v]. Query [CK78, HY79, Rei77]. queue [PS75]. Queueing [Mar79]. Queues [CP79]. Queuing [GM79, Spi79, TW79].

rams [Fra78]. Random

[GK78, PS75, Spi79]. Randomness [Knu77]. Ranges [Har77b]. Real [Alf77, Kod78a, Sch79, Mcc75a]. Real-Time [Alf77, Kod78a, Sch79, Mcc75a]. Reassignment [Bok79]. Recognition [Ano77t, Far77, SB77]. Records [LSM78]. Recoverability [ALS78]. recovery [CBDU75]. Recursive [CSA79]. Reduction [BJ76]. Redundant [BRGP78]. Referees [Ano76b, Ano77b, Ano78b, Ano79b]. references [GM75]. Refinement [Tow79]. Reflections [Ano77u]. Region [For78]. Relational [Rip77]. Reliability [Goe80, How76, LLS75, SW78, GH75, Mus75]. Reliable [SB78, Wul75, BMU75, Cul75]. Replacement [Wal77]. reply [Bak76]. Report [RS78]. Representation [Rip77]. Requirements [Alf77, Ano77u, BBD77, HKN79, Ott79, RS77]. Resolution [Wor76]. Resource [SB78, SKB77]. Resources [RL79]. Restructuring [GIL78]. Retrieval [SF77]. Review [Gli79]. Revocation [Gli79]. RIG [BFL^{+}76]. Road [PB78]. Rochester [BFL^{+}76]. rollback [CBDU75]. Room [CP79].

SA [Ros77]. Samples [BBS^{+}79]. scale [BMU75]. Scheduling [Bok79, Sto77a]. Schema [Rou79]. Science [BZ79, Hal79, OB79]. SDD [BRGP78]. SDD-1 [BRGP78]. SE [Ano76a, Ano77a, Ano78a, Ano79a]. SE-2 [Ano76a]. SE-3 [Ano77a]. SE-4 [Ano78a]. SE-5 [Ano79a]. Search [Ben79]. Searching [Gho76, LCC76]. Security [DLSW78]. Seem [Ano76m]. Selected [Ano78-29, Ano78-30]. Selecting [CHS77].

Selection [RT78, Spi79, GG75a, GG75b]. Sequence [CP79]. Sequences [Boo79, Mis78c]. Serializability [BSW79]. Serializers [HA79]. Servers [GM79]. Service [Spi79]. Set [Kat79, Smi78]. sets [Weg75]. Shapiro [LLS81]. Shared [JL76, RL79]. Simple [CH79]. Simulation [BJ76, CM79]. Simulator [HP78, Mcc75a]. Sizing [Put78]. Skeleton [CC78, CK78]. Small [DK76]. SMOTL [BBS^{+}79]. SNOBOL4 [RGH78]. Society [Ano76d, Ano76r, Ano76n, Ano76s, Ano76o, Ano76t, Ano76p, Ano76u, Ano76q, Ano76z, Ano77d, Ano77e, Ano77f, Ano77-28, Ano77v, Ano77-29, Ano77w, Ano77-30, Ano77x, Ano77-31, Ano77y, Ano77-32, Ano77z, Ano77-33, Ano77-27, Ano78g, Ano78-37, Ano78-38, Ano78-31, Ano78w, Ano78-32, Ano78x, Ano78-33, Ano78y, Ano78-29, Ano78-34, Ano78z, Ano78-35, Ano78-27, Ano78-36, Ano78-30, Ano78-28, Ano79f, Ano79g, Ano79h, Ano79-38, Ano79-29, Ano79w, Ano79-30, Ano79x, Ano79-31, Ano79y, Ano79-32, Ano79z, Ano79-33, Ano79-27, Ano79-34, Ano79-28]. Software [Ano76m, Ano76a, Ano76e, Ano76f, Ano76g, Ano76h, Ano77a, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78u, Ano78a, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n, BZ79, BBD77, BRH79, Boo79, CS78, Cho78, Coo78, CSM^{+}79, Dal77, DV77, DN78, GS78, Goe80, Hal79, HZ76, Har77a, HW77, HP78, IB77, LLS81, MW78b, Mer78, Mil76, OB79, Par79, PB78, Put78, RWS^{+}78, Rom77, SW78, SH79a, SH79b, Sha78, SB75, Wil76, BT75, BMU75, Chu75, Cul75, KY75, Mcc75a, Mus75, RKC75, RH75, Ran75, RDB75, Wul75]. Solution [Put78]. Solving [SO77]. Some [BMU75, Els76, LLS81, LLD79, Mis78b, Rob79]. Sorting [Sto78b]. source [Roc75]. Space [Chi78]. Special [Ano78u]. Specification [GR78, HS78, HA79, LZ75, PHW76, Spi76, Noo75]. Specifications [BGW78]. Standard [MS77, RW76]. Standing [Ano76d, Ano77d, Ano77e, Ano77f]. STAR [Sto78b]. State [Cho78]. Statement [Gri78]. Statements [Elg76a, Elg76b]. Status [Ano77c]. Still [Ano76m]. Storage [CC77, RGH78]. strategies [CBDU75]. Strategy [Har79, How76]. Streams [Ell78]. Strong [BY75]. STRUCT [SV75]. Structure [Ada77, BJ79, Cla79, Far77, GY76a, Zwe77, PS75, Ran75]. Structured [Bak75, Elg76a, Elg76b, Els77, IB77, Noo75, RS77, TH77, Tom76, Wil76, Mcc75b, RW76, Ros77]. Structures [Ano77t, Pra78, RT78, GM75, SB75]. Structuring [BS79]. Studies [How78b]. Study [CM79, CS79, LP79, Smi78, Wor76, Zwe77]. subscribe [Ano76v, Ano76w, Ano76x, Ano77-34, Ano77-35, Ano77-36, Ano77-37, Ano77-38, Ano79-36]. Subscription [Ano76d, Ano77d, Ano77e, Ano77f]. Subset [PHW76]. Suggests [Ano76m]. Summary [Ano78e, Goe80]. Support [DJ77]. Swapping [CC77]. Switching [And79]. Symbolic [CHT79, How77, How78a]. Symbolically [Cla76]. Synchronization [Sil79]. Synchronous [ABSS78]. Syndrome [Ros78]. Syntactic [Tai78]. Syntax [GR78, Led77]. Synthesis [CC78, MW79]. synthesizer [GW75]. System [And79, BFL^{+}76, BRGP78, BBS^{+}79, BRH79, Cer77, CC76, Cla76, DV77, Ell78, Far77, GT76, GIL78, HY79, How77, How78a, Nut77, PHW76, Ran75, SB77, WG76, Cul75, End75, GLB75, MBGK75, Roc75, SV75]. systematic [RKC75]. Systems [ABSS78, ALS78, Ano76k, BS76, BS79, Boo79, CHS77, Chi78, CC77, DJ77, DHI78, GS78, HP78, Kod78a, Kod78b, Led77, LLS81, LLD79, Sil79, Sto78a, SO77, TH77, CBDU75, RH75].

Table [Ano75p, Ano75q, Ano75r, Ano75s, Ano76e, Ano76f, Ano76g, Ano76h, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n]. Task [CS78]. Tasks [CSM^{+}79]. team [Voi75]. Technique [GY76a, Mis78c, TH77, BT75]. Techniques [HA79, Rob79, LZ75]. Technological [DN78]. Technology [Ano77c]. Test [Cla76, GMO76, MS76, RHC76, GG75a, GG75b]. Testing [Ano76m, Cho78, Ham77, How76, How77, How78a, How78b, Kun78, NH79, RH75, Sam77, Fai75, RKC75]. Text [WHH79]. Their [Smi78, End75, Pai75]. theorem [Moo75]. Theoretical [How78b]. theory [GG75a, GG75b, Mus75]. Third [Ano78u]. Throughput [CP79, Wal77]. Tidy [WS79]. Time [AC79, Alf77, Boc78, Cer77, Kod78a, Sch79, Mcc75a]. Timing [Wor76]. tolerance [Ran75]. Tolerant [GS78, LLS81]. Tone(R) [SF77]. Top [CH79, Mcc75b, Tow79]. Top-Down [CH79, Tow79, Mcc75b]. Total [Bab79]. Touch [SF77]. Touch-Tone(R) [SF77]. Toy [Ros78]. Trace [Smi77]. Transactions [Ano76a, Ano76e, Ano76f, Ano76g, Ano76h, Ano77a, Ano77g, Ano77h, Ano77i, Ano77j, Ano77k, Ano77l, Ano78f, Ano78v, Ano78a, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano79a, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n, Ano78f, Ano78v]. Transformation [CS79, Weg76]. Translation [CK78, CS79, Led77]. Tree [KSY77]. Trees [Ben79, WS79]. Tutorial [Ano78-29, Ano78-30, Ano79f, Ano79g, Ano79h, Ano76z]. Two [GMO76, Smi77, Sto78a]. Two-Processor [Sto78a]. Types [LP79].

uniform [GM75]. Unit [Woo79]. Units [Sto77b]. Unreliable [SB78]. Until [Ano76m]. Use [BZ79, Cla79, Ell78, Rei77, Smi78]. User [DHI78]. User-Perceived [DHI78]. Using [GR78, Kat78, LP79, Sto77b, SB75]. Utilization [Mas79].

V [HP78]. validation [RDB75]. Value [Har77b]. Variables [Har77b]. Verifiable [Ada77]. Verification [CM79, Mis78b, Pol79, RY76, Van79, Weg77, WLS76, BY75, GLB75]. Verifier [Mor79]. Versus [DK76]. Virtual [Ano76k, BS76, BS79, CC77, CH78, DJ77, GK78]. Vol [Ano76a, Ano77a, Ano78a, Ano79a].

Waiting [CP79]. Way [Mer78]. Weapons [MW78b]. Week79 [Ano79f, Ano79g, Ano79h]. well [Weg75]. well-founded [Weg75]. Wellmade [BP78]. Without [Elg76a, Elg76b].

X.25 [BJ79].

References

Akkoyunlu:1978:CES

[ABSS78] E. A. Akkoyunlu, A. J. Bernstein, F. B. Schneider, and A. Silberschatz. Conditions for the equivalence of synchronous and asynchronous systems. IEEE Transactions on Software Engineering, SE-4(6): 507–516, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                                  /ieeexplore.ieee.org/stamp/
                                                  stamp.jsp?arnumber=1702570.

Abrahams:1979:CTA

[AC79] P. W. Abrahams and L. A. Clarke. Compile-time analysis of data list-format list correspondences. IEEE Transactions on Software Engineering, SE-5(6):

612–617, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702678.

Adams:1977:GVI

[Ada77] J. M. Adams. A general, verifiable iterative control structure. IEEE Transactions on Software Engineering, SE-3(2):144–149, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702416.

Alford:1977:REM

[Alf77]  M. W. Alford.  A require-

ments engineering methodology for real-time processing requirements. IEEE Transactions on Software Engineering, SE-3(1):60–69, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702403.

Anderson:1978:MRM

[ALS78] T. Anderson, P. A. Lee, and S. K. Shrivastava. A model of recoverability in multilevel systems. IEEE Transactions on Software Engineering, SE-4(6): 486–494, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702568.

Andrews:1979:DMS

[And79] G. R. Andrews. The design of a message switching system: An application and evaluation of modula. IEEE Transactions on Software Engineering, SE-5(2):138–147, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702608.

Anonymous:1975:AI

[Ano75a] Anonymous. Author index. IEEE Transactions on Software Engineering, SE-1(4):1–3, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312877.

Anonymous:1975:BMa

[Ano75b] Anonymous. [back matter]. IEEE Transactions on Software Engineering, SE-1(1):1–4, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312828.

Anonymous:1975:BMb

[Ano75c] Anonymous. [back matter]. IEEE Transactions on Software Engineering, SE-1(2):1–5, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312846.

Anonymous:1975:BMc

[Ano75d] Anonymous. [back matter]. IEEE Transactions on Software Engineering, SE-1(3):1–4, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312859.

Anonymous:1975:BMd

[Ano75e] Anonymous. [back matter]. IEEE Transactions on Software Engineering, SE-1(4):1–5, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312878.

Anonymous:1975:ENa

[Ano75f] Anonymous. Editor’s notice.

IEEE Transactions on Software Engineering, SE-1(1):1–6, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312815.

Anonymous:1975:ENb

[Ano75g] Anonymous. Editor’s notice.

IEEE Transactions on Software Engineering, SE-1(2):137, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312832.

Anonymous:1975:ENc

[Ano75h] Anonymous. Editor’s notice.

IEEE Transactions on Software Engineering, SE-1(3):261, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6312850.

Anonymous:1975:FM

[Ano75i] Anonymous. [front matter]. IEEE Transactions on Software Engineering, SE-1(1):1–4, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312814.

Anonymous:1975:GE

[Ano75j] Anonymous. Guest editorial. IEEE Transactions on Software Engineering, SE-1(2):137–139, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312833.

Anonymous:1975:IA

[Ano75k] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-1(1):1, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312827.

Anonymous:1975:IFCa

[Ano75l] Anonymous. [inside front cover]. IEEE Transactions on Software Engineering, SE-1(1):c2, March 1975. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312813.

Anonymous:1975:IFCb

[Ano75m] Anonymous. [inside front cover].

IEEE Transactions on Software Engineering, SE-1(2):c2, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312831.

Anonymous:1975:IFCc

[Ano75n] Anonymous. [inside front cover].

IEEE Transactions on Software Engineering, SE-1(3):c2, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312849.

Anonymous:1975:IFCd

[Ano75o] Anonymous. [inside front cover].

IEEE Transactions on Software Engineering, SE-1(4):c2, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312862.

Anonymous:1975:TCa

[Ano75p] Anonymous. Table of contents. IEEE Transactions on Software Engineering, SE-1(1):1, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312812.

Anonymous:1975:TCb

[Ano75q] Anonymous. Table of contents. IEEE Transactions on Software Engineering, SE-1(2):1, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312830.

Anonymous:1975:TCc

[Ano75r] Anonymous. Table of contents. IEEE Transactions on Software Engineering, SE-1(3):1, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312848.

Anonymous:1975:TCd

[Ano75s] Anonymous. Table of contents. IEEE Transactions on Software Engineering, SE-1(4):1, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312861.

Anonymous:1976:IIT

[Ano76a] Anonymous. 1976 index IEEE Transactions on Software Engineering vol. SE-2. IEEE Transactions on Software Engineering, SE-2(4):329, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702391.

Anonymous:1976:AR

[Ano76b] Anonymous. Acknowledgment of referees. IEEE Transactions on Software Engineering, SE-2(4):329, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702390.

Anonymous:1976:AP

[Ano76c] Anonymous. Advance program. IEEE Transactions on Software Engineering, SE-2(3): 232, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702372.

Anonymous:1976:AIC

[Ano76d] Anonymous. Announcing the IEEE Computer Society’s standing subscription plan [advertisement]. IEEE Transactions on Software Engineering, SE-2(2): 137, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702355.

Anonymous:1976:ITSa

[Ano76e] Anonymous. IEEE Transactions on Software Engineering

— table of contents. IEEE Transactions on Software Engineering, SE-2(1):c1, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702330.

Anonymous:1976:ITSb

[Ano76f] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-2(2): c1, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702342.

Anonymous:1976:ITSc

[Ano76g] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-2(3):c1, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702359.

Anonymous:1976:ITSd

[Ano76h] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-2(4):c1, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702376.

Anonymous:1976:CP

[Ano76i] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-2(3): 232, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702375.

Anonymous:1976:CMM

[Ano76j] Anonymous. Computers … by the millions, for the millions. IEEE Transactions on Software Engineering, SE-2(2):137–e–137– e, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702358.

Anonymous:1976:C

[Ano76k] Anonymous. Correction to “Dynamic Improvement of Locality in Virtual Memory Systems”. IEEE Transactions on Software Engineering, SE-2(2): 137, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702353.

See [BS76].

Anonymous:1976:GE

[Ano76l] Anonymous. Guest editorial.

IEEE Transactions on Software Engineering, SE-2(4):237, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702378.

Anonymous:1976:GEC

[Ano76m] Anonymous. Guest editorial: A case for software testing or why everything we seem to learn about software development still suggests the need for testing or what to do until we achieve “perfection”. IEEE Transactions on Software Engineering, SE-2(3):194, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702365.

Anonymous:1976:ICSb

[Ano76n] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-2(1):c2, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702331.

Anonymous:1976:ICSd

[Ano76o] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-2(2):c2, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702343.

Anonymous:1976:ICSf

[Ano76p] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-2(3): c2, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702360.

Anonymous:1976:ICSh

[Ano76q] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-2(4):c2, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702377.

Anonymous:1976:ICSa

[Ano76r] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-2(1):68–a–68–a, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702340.

Anonymous:1976:ICSc

[Ano76s] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-2(2):137–d–137–d, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702357.

Anonymous:1976:ICSe

[Ano76t] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-2(3):232–e–232–e, September 1976. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702374.

Anonymous:1976:ICSg

[Ano76u] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-2(4):329–d–329–d, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702392.

Anonymous:1976:IYCa

[Ano76v] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Software Engineering, SE-2(2):137, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702356.

Anonymous:1976:IYCb

[Ano76w] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Software Engineering, SE-2(3):232, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702373.

Anonymous:1976:IYCc

[Ano76x] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Software Engineering, SE-2(4):329, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702393.

Anonymous:1976:MAC

[Ano76y] Anonymous. Management auditing of computer operations. IEEE Transactions on Software Engineering, SE-2(1):68–b–68–b, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702341.

Anonymous:1976:OTY

[Ano76z] Anonymous. Organize a tutorial for your own IEEE Computer Society chapter. IEEE Transactions on Software Engineering, SE-2(2):137–a–137–a, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702354.

Anonymous:1977:IIT

[Ano77a] Anonymous. 1977 index IEEE Transactions on Software Engineering vol. SE-3. IEEE Transactions on Software Engineering, SE-3(6):480, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702478.

Anonymous:1977:AR

[Ano77b] Anonymous. Acknowledgment of referees. IEEE Transactions on Software Engineering, SE-3(1):102, January/February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702407.

Anonymous:1977:ACT

[Ano77c] Anonymous. Announcing computer technology: Status, limits, alternatives. IEEE Transactions on Software Engineering, SE-3(6):480, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702481.

Anonymous:1977:AICa

[Ano77d] Anonymous. Announcing the IEEE Computer Society’s standing subscription plan [advertisement]. IEEE Transactions on Software Engineering, SE-3(1):102, January/February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702411.

Anonymous:1977:AICb

[Ano77e] Anonymous. Announcing the IEEE Computer Society’s standing subscription plan [advertisement]. IEEE Transactions on Software Engineering, SE-3(4):330, July/August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702452.

Anonymous:1977:AICc

[Ano77f] Anonymous. Announcing the IEEE Computer Society’s standing subscription plan [advertisement]. IEEE Transactions on Software Engineering, SE-3(5):374, September/ October 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702464.

Anonymous:1977:ITSa

[Ano77g] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-3(1):c1, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702394.

Anonymous:1977:ITSb

[Ano77h] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-3(2):c1, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702412.

Anonymous:1977:ITSc

[Ano77i] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-3(3):c1, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702426.

Anonymous:1977:ITSd

[Ano77j] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-3(4):c1, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702440.

Anonymous:1977:ITSe

[Ano77k] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-3(5):c1, September/ October 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702455.

Anonymous:1977:ITSf

[Ano77l] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-3(6):c1, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702467.

Anonymous:1977:CPa

[Ano77m] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-3(2):192, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702425.

Anonymous:1977:CPb

[Ano77n] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-3(3):262, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702436.

Anonymous:1977:CPc

[Ano77o] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-3(4):330, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702454.

Anonymous:1977:CPd

[Ano77p] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-3(5):374, September/October 1977. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702466.

Anonymous:1977:CF

[Ano77q] Anonymous. Compcon77 Fall. IEEE Transactions on Software Engineering, SE-3(3):262, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702439.

Anonymous:1977:ENa

[Ano77r] Anonymous. Editor’s notice. IEEE Transactions on Software Engineering, SE-3(1):1, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702396.

Anonymous:1977:ENb

[Ano77s] Anonymous. Editor’s notice. IEEE Transactions on Software Engineering, SE-3(4):265, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702442.

Anonymous:1977:GED

[Ano77t] Anonymous. Guest editorial: Data structures and pattern recognition. IEEE Transactions on Software Engineering, SE-3(2):160–161, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702419.

Anonymous:1977:GER

[Ano77u] Anonymous. Guest editorial: Reflections on requirements. IEEE Transactions on Software Engineering, SE-3(1):2–5, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702397.

Anonymous:1977:ICSb

[Ano77v] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-3(1):c2, January/February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702395.

Anonymous:1977:ICSd

[Ano77w] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-3(2):c2, March/April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702413.

Anonymous:1977:ICSf

[Ano77x] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-3(3): c2, May/June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702427.

Anonymous:1977:ICSh

[Ano77y] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-3(4):c2, July/August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702441.

Anonymous:1977:ICSj

[Ano77z] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-3(5):c2, September/ October 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702456.

Anonymous:1977:ICSl

[Ano77-27] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-3(6):c2, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702468.

Anonymous:1977:ICSa

[Ano77-28] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-3(1):102–c–102–c, January/ February 1977. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702410.

Anonymous:1977:ICSc

[Ano77-29] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-3(2):192–a–192–a, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702424.

Anonymous:1977:ICSe

[Ano77-30] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-3(3):262–c–262–c, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702438.

Anonymous:1977:ICSg

[Ano77-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-3(4):330–c–330–c, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702453.

Anonymous:1977:ICSi

[Ano77-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-3(5):374–c–374–c, September/ October 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702465.

Anonymous:1977:ICSk

[Ano77-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-3(6):480–e–480–e, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702480.

Anonymous:1977:IYCa

[Ano77-34] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Software Engineering, SE-3(1):102, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702409.

Anonymous:1977:IYCb

[Ano77-35] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Software Engineering, SE-3(3):262, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702437.

Anonymous:1977:IYCc

[Ano77-36] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Software Engineering, SE-3(4):330, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702451.

Anonymous:1977:IYCd

[Ano77-37] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Software Engineering, SE-3(5):374, September/ October 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702463.

Anonymous:1977:IYCe

[Ano77-38] Anonymous. If your computer engineering library doesn’t subscribe to all 3 it’s not complete. IEEE Transactions on Software Engineering, SE-3(6):480, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702479.

Anonymous:1977:IA

[Ano77-39] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-3(1):102, January/February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702408.

Anonymous:1978:IIT

[Ano78a] Anonymous. 1978 index IEEE Transactions on Software Engineering vol. SE-4. IEEE Transactions on Software Engineering, SE-4(6):533, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702575.

Anonymous:1978:AR

[Ano78b] Anonymous. Acknowledgment of referees. IEEE Transactions on Software Engineering, SE-4(1):78, January/February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702495.

Anonymous:1978:AP

[Ano78c] Anonymous. Advance program. IEEE Transactions on Software Engineering, SE-4(2):148, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702510.

Anonymous:1978:APA

[Ano78d] Anonymous. Advance program announcement. IEEE Transactions on Software Engineering, SE-4(3):259, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702529.

Anonymous:1978:APS

[Ano78e] Anonymous. Advance program summary. IEEE Transactions on Software Engineering, SE-4(4):361, July/August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702545.

Anonymous:1978:ANQ

[Ano78f] Anonymous. Announcing a new quarterly IEEE Transactions … IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Software Engineering, SE-4(5):448, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702561.

Anonymous:1978:BCP

[Ano78g] Anonymous. Books for computer professionals from the IEEE Computer Society new publications. IEEE Transactions on Software Engineering, SE-4(1):78–a–78–b, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702496.

Anonymous:1978:ITSa

[Ano78h] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-4(1):c1, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702482.

Anonymous:1978:ITSb

[Ano78i] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-4(2):c1, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702499.

Anonymous:1978:ITSc

[Ano78j] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-4(3):c1, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702514.

Anonymous:1978:ITSd

[Ano78k] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-4(4):c1, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702530.

Anonymous:1978:ITSe

[Ano78l] Anonymous. IEEE Transactions on Software Engineering

— table of contents. IEEE Transactions on Software Engineering, SE-4(5):c1, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702549.

Anonymous:1978:ITSf

[Ano78m] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-4(6):c1, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702562.

Anonymous:1978:CPa

[Ano78n] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-4(1):78, January/February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702498.

Anonymous:1978:CPb

[Ano78o] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-4(2):148, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702511.

Anonymous:1978:CPc

[Ano78p] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-4(2):148, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702509.

Anonymous:1978:CPd

[Ano78q] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-4(4):361, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702548.

Anonymous:1978:CPe

[Ano78r] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-4(6):533, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702579.

Anonymous:1978:EN

[Ano78s] Anonymous. Editor’s notice. IEEE Transactions on Software Engineering, SE-4(1):1–2, January/February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702484.

Anonymous:1978:GE

[Ano78t] Anonymous. Guest editorial.

IEEE Transactions on Software Engineering, SE-4(3):153, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702516.

Anonymous:1978:GES

[Ano78u] Anonymous. Guest editorial for the special collection from the Third International Conference on Software Engineering. IEEE Transactions on Software Engineering, SE-4(4):261–262, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702532.

Anonymous:1978:IAN

[Ano78v] Anonymous. IEEE announcing a new quarterly IEEE Transactions … IEEE Transactions on Pattern Analysis and Machine Intelligence. IEEE Transactions on Software Engineering, SE-4(6):533–g–533–g, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702577.

Anonymous:1978:ICSb

[Ano78w] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-4(1):c2, January/February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702483.

Anonymous:1978:ICSd

[Ano78x] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-4(2):c2, March/April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702500.

Anonymous:1978:ICSf

[Ano78y] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-4(3): c2, May/June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702515.

Anonymous:1978:ICSi

[Ano78z] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-4(4):c2, July/August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702531.

Anonymous:1978:ICSk

[Ano78-27] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-4(5):c2, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702550.

Anonymous:1978:ICSn

[Ano78-28] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-4(6):c2, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702563.

Anonymous:1978:ICSg

[Ano78-29] Anonymous. The IEEE Computer Society announces tutorial and selected papers in digital image processing. IEEE Transactions on Software Engineering, SE-4(4):361–c–361–c, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702546.

Anonymous:1978:ICSm

[Ano78-30] Anonymous. The IEEE Computer Society announces tutorial and selected papers in digital image processing. IEEE Transactions on Software Engineering, SE-4(6):533–f–533–f, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702576.

Anonymous:1978:ICSa

[Ano78-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-4(1):78–c–78–c, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702497.

Anonymous:1978:ICSc

[Ano78-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-4(2):148–e–148–e, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702512.

Anonymous:1978:ICSe

[Ano78-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-4(3):259–b–259–b, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702528.

Anonymous:1978:ICSh

[Ano78-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-4(4):361–d–361–d, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702547.

Anonymous:1978:ICSj

[Ano78-35] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-4(5):448–a–448–a, September/October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702560.

Anonymous:1978:ICSl

[Ano78-36] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-4(6):533–h–533–h, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702578.

Anonymous:1978:IASa

[Ano78-37] Anonymous. Information about the Society. IEEE Transactions on Software Engineering, SE-4(2):148–f–148–f, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702513.

Anonymous:1978:IASb

[Ano78-38] Anonymous. Information about the Society. IEEE Transactions on Software Engineering, SE-4(3):259–a–259–a, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702527.

Anonymous:1978:IA

[Ano78-39] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-4(3): 259, May/June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702526.

Anonymous:1979:IIT

[Ano79a] Anonymous. 1979 index IEEE Transactions on Software Engineering vol. SE-5. IEEE Transactions on Software Engineering, SE-5(6):659, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702683.

Anonymous:1979:AR

[Ano79b] Anonymous. Acknowledgment of referees. IEEE Transactions on Software Engineering, SE-5(1):67–68, January/ February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702589.

Anonymous:1979:AA

[Ano79c] Anonymous. Advance announcement. IEEE Transactions on Software Engineering, SE-5(2):174, March/April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702612.

Anonymous:1979:APAa

[Ano79d] Anonymous. Advance program announcement. IEEE Transactions on Software Engineering, SE-5(1):71, January/ February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702594.

Anonymous:1979:APAb

[Ano79e] Anonymous. Advance program announcement. IEEE Transactions on Software Engineering, SE-5(2):174, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702614.

Anonymous:1979:AICa

[Ano79f] Anonymous. Announcing … the IEEE Computer Society’s tutorial week79. IEEE Transactions on Software Engineering, SE-5(3):289–e–289–e, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702632.

Anonymous:1979:AICb

[Ano79g] Anonymous. Announcing … the IEEE Computer Society’s tutorial week79. IEEE Transactions on Software Engineering, SE-5(4):428–b–428–b, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702649.

Anonymous:1979:AICc

[Ano79h] Anonymous. Announcing … the IEEE Computer Society’s tutorial week79. IEEE Transactions on Software Engineering, SE-5(5):538–d–538–d, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702667.

Anonymous:1979:ITSa

[Ano79i] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-5(1):c1, January/ February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702580.

Anonymous:1979:ITSb

[Ano79j] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-5(2):c1, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702595.

Anonymous:1979:ITSc

[Ano79k] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-5(3):c1, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702615.

Anonymous:1979:ITSd

[Ano79l] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-5(4):c1, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702633.

Anonymous:1979:ITSe

[Ano79m] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-5(5):c1, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702650.

Anonymous:1979:ITSf

[Ano79n] Anonymous. IEEE Transactions on Software Engineering — table of contents. IEEE Transactions on Software Engineering, SE-5(6):c1, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702668.

Anonymous:1979:CPa

[Ano79o] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-5(1):71, January/February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702592.

Anonymous:1979:CPb

[Ano79p] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-5(2):174, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702611.

Anonymous:1979:CPc

[Ano79q] Anonymous. Call for papers. IEEE Transactions on Software Engineering, SE-5(3):289, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702628.

Anonymous:1979:CPd

[Ano79r] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-5(5):538, September/October 1979. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702665.

Anonymous:1979:CPe

[Ano79s] Anonymous. Call for papers.

IEEE Transactions on Software Engineering, SE-5(6):659, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702685.

Anonymous:1979:ENc

[Ano79t] Anonymous. Editor’s note. IEEE Transactions on Software Engineering, SE-5(6):541, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702670.

Anonymous:1979:ENa

[Ano79u] Anonymous. Editor’s notice.

IEEE Transactions on Software Engineering, SE-5(2):73, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702597.

Anonymous:1979:ENb

[Ano79v] Anonymous. Editor’s notice. IEEE Transactions on Software Engineering, SE-5(4):293, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702635.

Anonymous:1979:ICSb

[Ano79w] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-5(1):c2, January/February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702581.

Anonymous:1979:ICSd

[Ano79x] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-5(2):c2, March/April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702596.

Anonymous:1979:ICSf

[Ano79y] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-5(3): c2, May/June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702616.

Anonymous:1979:ICSh

[Ano79z] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-5(4):c2, July/August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702634.

Anonymous:1979:ICSj

[Ano79-27] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-5(5):c2, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702651.

Anonymous:1979:ICSl

[Ano79-28] Anonymous. IEEE Computer Society. IEEE Transactions on Software Engineering, SE-5(6):c2, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702669.

Anonymous:1979:ICSa

[Ano79-29] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-5(1):71–b–71–b, January/ February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702593.

Anonymous:1979:ICSc

[Ano79-30] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-5(2):174–c–174–c, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702613.

Anonymous:1979:ICSe

[Ano79-31] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-5(3):289–d–289–d, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702631.

Anonymous:1979:ICSg

[Ano79-32] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-5(4):428–a–428–a, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702648.

Anonymous:1979:ICSi

[Ano79-33] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-5(5):538–c–538–c, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702666.

Anonymous:1979:ICSk

[Ano79-34] Anonymous. IEEE Computer Society publications. IEEE Transactions on Software Engineering, SE-5(6):659–f–659–f, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702684.

Anonymous:1979:ICF

[Ano79-35] Anonymous. IEEE copyright form. IEEE Transactions on Software Engineering, SE-5(1):69–70, January/February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702590.

Anonymous:1979:IYC

[Ano79-36] Anonymous. if your computer engineering library doesn’t subscribe to all 4 it’s not complete. IEEE Transactions on Software Engineering, SE-5(3):289, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702629.

Anonymous:1979:MMH

[Ano79-37] Anonymous. In memory of Maurice H. Halstead. IEEE Transactions on Software Engineering, SE-5(2):74, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702598.

Anonymous:1979:IAS

[Ano79-38] Anonymous. Information about the Society. IEEE Transactions on Software Engineering, SE-5(3):289–c–289–c, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702630.

Anonymous:1979:IA

[Ano79-39] Anonymous. Information for authors. IEEE Transactions on Software Engineering, SE-5(1):71, January/February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702591.

Babich:1979:PTC

[Bab79] A. F. Babich. Proving total correctness of parallel programs. IEEE Transactions on Software Engineering, SE-5(6): 558–574, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702673.

Baker:1975:SPP

[Bak75] F. T. Baker. Structured programming in a production programming environment. IEEE Transactions on Software Engineering, SE-1(2):241–252, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312844.

See comments [Tom76].

Baker:1976:ARS

[Bak76] F. T. Baker. Author’s reply. IEEE Transactions on Software Engineering, SE-2(1):67–68, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702339.

Bell:1977:EAC

[BBD77] T. E. Bell, D. C. Bixler, and M. E.

Dyer. An extendable approach to computer-aided software requirements engineering. IEEE Transactions on Software Engineering, SE-3(1):49–60, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702402.

Bicevskis:1979:SSC

[BBS^{+}79] J. Bicevskis, J. Borzovs, U. Straujums, A. Zarins, and E. F. Miller. SMOTL — a system to construct samples for data processing program debugging. IEEE Transactions on Software Engineering, SE-5(1):60–66, January/ February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702588.

Baer:1977:MDE

[BE77] J. L. Baer and C. S. Ellis. Model, design, and evaluation of a compiler for a parallel processing environment. IEEE Transactions on Software Engineering, SE-3(6): 394–405, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702471.

Bentley:1979:MBS

[Ben79] J. L. Bentley. Multidimensional binary search trees in database applications. IEEE Transactions on Software Engineering, SE-5(4):333–340, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702638.

Ball:1976:RRI

[BFL^{+}76] J. E. Ball, J. Feldman, J. R. Low, R. Rashid, and P. Rovner. RIG, Rochester’s Intelligent gateway: System overview. IEEE Transactions on Software Engineering, SE-2(4):321–328, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702389.

Balzer:1978:IPS

[BGW78] R. Balzer, N. Goldman, and D. Wile. Informality in program specifications. IEEE Transactions on Software Engineering, SE-4(2):94–103, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702503.

Birman:1976:PRA

[BJ76] A. Birman and W. H. Joyner.

A problem-reduction approach to proving simulation between programs. IEEE Transactions on Software Engineering, SE-2 (2):87–96, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702346.

Bochmann:1979:DSX

[BJ79] G. V. Bochmann and T. Joachim.

Development and structure of an X.25 implementation. IEEE Transactions on Software Engineering, SE-5(5):429–439, September/October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702652.

Biermann:1976:CPE

[BK76] A. W. Biermann and R. Krishnaswamy. Constructing programs from example computations. IEEE Transactions on Software Engineering, SE-2(3):141– 153, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702361.

Basu:1975:PLP

[BM75] S. K. Basu and J. Misra. Proving loop programs. IEEE Transactions on Software Engineering, SE-1(1):76–86, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312822.

Boehm:1975:SEA

[BMU75] B. W. Boehm, R. K. Mcclean, and D. E. Urfrig. Some experience with automated aids to the design of large-scale reliable software. IEEE Transactions on Software Engineering, SE-1(1):125–133, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312826.

Bochmann:1978:CTM

[Boc78] G. V. Bochmann. Compile time memory allocation for parallel processes. IEEE Transactions on Software Engineering, SE-4(6): 517–520, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702571.

Bokhari:1979:DPS

[Bok79] S. H. Bokhari. Dual processor scheduling with dynamic reassignment. IEEE Transactions on Software Engineering, SE-5(4):341–349, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702639.

Booth:1979:POS

[Boo79] T. L. Booth. Performance optimization of software systems processing information sequences modeled by probabilistic languages. IEEE Transactions on Software Engineering, SE-5(1):31–44, January/ February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702585.

Boyd:1978:IWD

[BP78] D. L. Boyd and A. Pizzarello.

Introduction to the wellmade design methodology. IEEE Transactions on Software Engineering, SE-4(4):276–282, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702535.

Bernstein:1978:CCM

[BRGP78] P. A. Bernstein, J. B. Rothnie, N. Goodman, and C. A. Papadimitriou. The concurrency control mechanism of SDD-1: A system for distributed databases (the fully redundant case). IEEE Transactions on Software Engineering, SE-4(3):154–168, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702517.

Boisvert:1979:SPE

[BRH79] Ronald F. Boisvert, John R. Rice, and Elias N. Houstis. A system for performance evaluation of partial differential equations software. IEEE Transactions on Software Engineering, SE-5(4):418–425, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702646.

BrinchHansen:1975:PLC

[Bri75] Per Brinch Hansen. The programming language Concurrent Pascal. IEEE Transactions on Software Engineering, SE-1(2):199– 207, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312840.

Baier:1976:DIL

[BS76] J. L. Baier and G. R. Sager. Dynamic improvement of locality in virtual memory systems. IEEE Transactions on Software Engineering, SE-2(1):54–62, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702336.

See correction [Ano76k].

Barrese:1979:SPE

[BS79] A. L. Barrese and S. D. Shariro.

Structuring programs for efficient operation in virtual memory systems. IEEE Transactions on Software Engineering, SE-5(6): 643–652, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702681.

Bernstein:1979:FAS

[BSW79] P. A. Bernstein, D. W. Shipman, and W. S. Wong. Formal aspects of serializability in database concurrency control. IEEE Transactions on Software Engineering, SE-5(3):203–216, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702620.

Basil:1975:IEP

[BT75] V. R. Basil and A. J. Turner.

Iterative enhancement: A practical technique for software development. IEEE Transactions on Software Engineering, SE-1(4):390–396, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312870.

Basu:1975:SVP

[BY75] S. K. Basu and R. T. Yeh.

Strong verification of programs.

IEEE Transactions on Software Engineering, SE-1(3):339– 346, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6312858.

Baker:1979:USS

[BZ79] A. L. Baker and S. H. Zweben. The use of software science in evaluating modularity concepts. IEEE Transactions on Software Engineering, SE-5(2):110–120, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702605.

Chandy:1975:AMR

[CBDU75] K. M. Chandy, J. C. Browne, C. W. Dissly, and W. R. Uhrig. Analytic models for rollback and recovery strategies in data base systems. IEEE Transactions on Software Engineering, SE-1(1):100–110, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312824.

Chu:1976:IHL

[CC76] Yaohan Chu and R. Cannon. Interactive high-level language direct-execution microprocessor system. IEEE Transactions on Software Engineering, SE-2(2):126–134, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702351.

Chow:1977:ASP

[CC77] We-Min Chow and W. W. Chiu.

An analysis of swapping policies in virtual storage systems. IEEE Transactions on Software Engineering, SE-3(2):150–156, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702417.

Chang:1978:DSAa

[CC78] Shi-Kuo Chang and Wu-Haung Cheng. Database skeleton and its application to logical database synthesis. IEEE Transactions on Software Engineering, SE-4(1):18–30, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702486.

Cernicica:1977:EAA

[Cer77] P. D. Cernicica. Evaluation of the average access time for a digital cassette memory system. IEEE Transactions on Software Engineering, SE-3(4):310–324, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702449.

Cremers:1978:FMV

[CH78] A. B. Cremers and T. N. Hibbard. Formal modeling of virtual machines. IEEE Transactions on Software Engineering, SE-4(5):426–436, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702557.

Comer:1979:SET

[CH79] D. Comer and M. H. Halstead. A simple experiment in topdown design. IEEE Transactions on Software Engineering, SE-5(2):105–109, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702604.

Chen:1978:PCP

[Che78] E. T. Chen. Program complexity and programmer productivity. IEEE Transactions on Software Engineering, SE-4(3):187–194, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702520.

Chin:1978:A

[Chi78] Y. H. Chin. An analysis of “distributed free space” in an operating and data management systems environment. IEEE Transactions on Software Engineering, SE-4(5):436–440, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702558.

Chow:1978:TSD

[Cho78] T. S. Chow. Testing software design modeled by finitestate machines. IEEE Transactions on Software Engineering, SE-4(3):178–187, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702519.

Chandy:1977:SCC

[CHS77] K. M. Chandy, J. Hogarth, and C. H. Sauer. Selecting capacities in computer communication systems. IEEE Transactions on Software Engineering, SE-3(4):290–295, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702445.

Cheatham:1979:SEA

[CHT79] T. E. Cheatham, G. H. Holloway, and J. A. Townley. Symbolic evaluation and the analysis of programs. IEEE Transactions on Software Engineering, SE-5(4):402–417, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702645.

Chu:1975:MSE

[Chu75] Y. Chu. A methodology for software engineering. IEEE Transactions on Software Engineering, SE-1(3):262–270, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312851.

Chang:1978:DSAb

[CK78] Shi-Kuo Chang and Jyh-Sheng Ke. Database skeleton and its application to fuzzy query translation. IEEE Transactions on Software Engineering, SE-4(1):31–44, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702487.

Camp:1977:IPN

[CL77] W. V. Camp and T. G. Lewis. Implementing a pseudorandom number generator on a minicomputer. IEEE Transactions on Software Engineering, SE-3(3):259–262, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702434.

Clarke:1976:SGT

[Cla76] L. A. Clarke. A system to generate test data and symbolically execute programs. IEEE Transactions on Software Engineering, SE-2(3):215–222, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702368.

Clark:1979:MDL

[Cla79] D. W. Clark. Measurements of dynamic list structure use in Lisp. IEEE Transactions on Software Engineering, SE-5(1):51–59, January/February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702587.

Chandy:1979:DSC

[CM79] K. M. Chandy and J. Misra. Distributed simulation: A case study in design and verification of distributed programs. IEEE Transactions on Software Engineering, SE-5(5):440–452, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702653.

Cooper:1978:CLS

[Coo78] J. D. Cooper. Corporate level software management. IEEE Transactions on Software Engineering, SE-4(4):319–326, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702541.

Caseau:1979:TCS

[CP79] P. Caseau and G. Pujolle. Throughput capacity of a sequence of queues with blocking due to finite waiting room. IEEE Transactions on Software Engineering, SE-5(6):631–642, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702680.

Cave:1978:CSL

[CS78] William C. Cave and Alan B. Salisbury. Controlling the software life cycle — the project management task. IEEE Transactions on Software Engineering, SE-4(4):326–334, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702542.

Cohen:1979:CSP

[CS79] J. Cohen and R. Sitver. A case study in program transformation: Translation into Polish. IEEE Transactions on Software Engineering, SE-5(6): 593–606, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702676.

Cohen:1979:EIR

[CSA79] J. Cohen, R. Sitver, and D. Auty. Evaluating and improving recursive descent parsers. IEEE Transactions on Software Engineering, SE-5(5):472–480, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702657.

Curtis:1979:MPC

[CSM^{+}79] B. Curtis, S. B. Sheppard, P. Milliman, M. A. Borst, and T. Love. Measuring the psychological complexity of software maintenance tasks with the Halstead and McCabe metrics. IEEE Transactions on Software Engineering, SE-5(2):96–104, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702603.

Culpepper:1975:SRE

[Cul75]  L. M. Culpepper.  A sys-

tem for reliable engineering software. IEEE Transactions on Software Engineering, SE-1(2):174– 178, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312837.

Daly:1977:MSD

[Dal77] E. B. Daly. Management of software development. IEEE Transactions on Software Engineering, SE-3(3):229–242, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702431.

Demillo:1978:EDB

[DDL78] R. A. Demillo, D. Dobkin, and R. J. Lipton. Even data bases that lie can be compromised. IEEE Transactions on Software Engineering, SE-4(1):73–75, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702492.

Dzida:1978:UPQ

[DHI78] W. Dzida, S. Herda, and W. D. Itzfeldt. User-perceived quality of interactive systems. IEEE Transactions on Software Engineering, SE-4(4):270–276, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702534.

Donovan:1977:VMC

[DJ77] J. J. Donovan and H. D. Jacqby. Virtual machine communication for the implementation of decision support systems. IEEE Transactions on Software Engineering, SE-3(5):333–342, September/ October 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702457.

DeRemer:1976:PLV

[DK76] F. DeRemer and H. H. Kron. Programming-in-the-large versus programming-in-the-small. IEEE Transactions on Software Engineering, SE-2(2):80–86, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702345.

Davida:1978:DS

[DLSW78] G. I. Davida, D. J. Linton, C. R.

Szelag, and D. L. Wells. Database security. IEEE Transactions on Software Engineering, SE-4(6): 531–533, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702574.

Dershowitz:1977:EPA

[DM77] N. Dershowitz and Z. Manna.

The evolution of programs: Automatic program modification. IEEE Transactions on Software Engineering, SE-3(6): 377–385, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702469.

DeRoze:1978:SLC

[DN78] B. C. De Roze and T. H. Nyman. The software life cycle — a management and technological challenge in the Department of Defense. IEEE Transactions on Software Engineering, SE-4(4):309–318, July/ August 1978. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702540.

Davis:1977:SDS

[DV77] C. G. Davis and C. R. Vick. The software development system. IEEE Transactions on Software Engineering, SE-3(1):69–84, January/February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702404.

Ekanadham:1979:CC

[EB79] K. Ekanadham and A. J. Bernstein. Conditional capabilities. IEEE Transactions on Software Engineering, SE-5 (5):458–464, September/October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702655.

Elgot:1976:EC

[Elg76a] C. C. Elgot. Erratum and corrigendum for “Structured Programming With and Without GO TO Statements”. IEEE Transactions on Software Engineering, SE-2(3):232, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702371.

See [Elg76b].

Elgot:1976:SPG

[Elg76b] C. C. Elgot. Structured programming with and without GO TO statements. IEEE Transactions on Software Engineering, SE-2(1):41–54, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702335.

See erratum and corrigendum [Elg76a].

Ellis:1978:IUM

[Ell78]  R. A. Ellis.  On the inter-

active use of a macroprocessor to generate operating system batch streams. IEEE Transactions on Software Engineering, SE-4(2):146–148, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702508.

Elshoff:1976:ASC

[Els76] J. L. Elshoff. An analysis of some commercial PL/I programs. IEEE Transactions on Software Engineering, SE-2(2):113–120, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702349.

Elshoff:1977:ISP

[Els77] J. L. Elshoff. The influence of structured programming on PL/I program profiles. IEEE Transactions on Software Engineering, SE-3(5):364–368, September/ October 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702461.

Endres:1975:AET

[End75] A. Endres. An analysis of errors and their causes in system programs. IEEE Transactions on Software Engineering, SE-1(2): 140–149, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6312834.

Fairley:1975:EPT

[Fai75] R. E. Fairley. An experimental program-testing facility. IEEE Transactions on Software Engineering, SE-1(4):350–357, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312864.

Fairley:1979:AAL

[Fai79] R. E. Fairley. Aladdin: Assembly language assertion driven debugging interpreter. IEEE Transactions on Software Engineering, SE-5(4):426–428, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702647.

Farag:1977:IPR

[Far77] R. F. H. Farag. Interactive pattern recognition: A system and data structure. IEEE Transactions on Software Engineering, SE-3(2):162–168, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702420.

Ford:1978:IGC

[For78] W. S. Ford. Implementation of a generalized critical region construct. IEEE Transactions on Software Engineering, SE-4(6): 449–455, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702564.

Francez:1978:AMA

[Fra78]  N. Francez.  An application

of a method for analysis of cyclic prog rams. IEEE Transactions on Software Engineering, SE-4(5):371–378, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702552.

Gait:1977:NNP

[Gai77]  J. Gait.  A new nonlin-

ear pseudorandom number generator. IEEE Transactions on Software Engineering, SE-3 (5):359–363, September/October 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702460.

Goodenough:1975:CTT

[GG75a] J. B. Goodenough and S. L. Gerhart. Correction to “Toward a theory of test data selection”. IEEE Transactions on Software Engineering, SE-1(4):425–426, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312876.

See [GG75b].

Goodenough:1975:TTT

[GG75b] J. B. Goodenough and S. L. Gerhart. Toward a theory of test data selection. IEEE Transactions on Software Engineering, SE-1(2):156–173, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312836.

See correction [GG75a].

Gannon:1975:LDP

[GH75] J. D. Gannon and J. J. Horning. Language design for programming reliability. IEEE Transactions on Software Engineering, SE-1(2):179–191, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312838.

Ghosh:1976:DDB

[Gho76] S. P. Ghosh. Distributing a data base with logical associations on a computer network for parallel searching. IEEE Transactions on Software Engineering, SE-2(2):106–113, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702348.

Goullon:1978:DRE

[GIL78] H. Goullon, R. Isle, and K. P.

Lohr. Dynamic restructuring in an experimental operating system. IEEE Transactions on Software Engineering, SE-4(4):298–307, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702538.

Gelenbe:1978:RIC

[GK78] E. Gelenbe and A. Kurinckx. Random injection control of multiprogramming in virtual memory. IEEE Transactions on Software Engineering, SE-4(1):2–17, January/February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702485.

Gligor:1979:OMA

[GL79] V. D. Gligor and B. G. Lindsay. Object migration and authentication. IEEE Transactions on Software Engineering, SE-5(6): 607–611, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702677.

Good:1975:IPV

[GLB75] D. I. Good, R. L. London, and W. W. Bledsoe. An interactive program verification system. IEEE Transactions on Software Engineering, SE-1(1):59–67, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312820.

Gligor:1979:RRA

[Gli79]  V. D. Gligor.  Review and

revocation of access privileges distributed through capabilities. IEEE Transactions on Software Engineering, SE-5(6): 575–586, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702674.

Geschke:1975:PUR

[GM75] C. M. Geschke and J. G. Mitchell. On the problem of uniform references to data structures. IEEE Transactions on Software Engineering, SE-1(2):207– 219, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312841.

Gonnet:1979:ACQ

[GM79] G. H. Gonnet and D. E.

Morgan. Analysis of closed queuing networks with periodic servers. IEEE Transactions on Software Engineering, SE-5(6): 653–659, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702682.

Gabow:1976:TPG

[GMO76] H. N. Gabow, S. N. Maheshwari, and L. J. Osterweil. On two problems in the generation of program test paths. IEEE Transactions on Software Engineering, SE-2(3):227–231, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702370.

Goel:1980:SD

[Goe80] A. L. Goel. A summary of the discussion on “An Analysis of Competing Software Reliability Models”. IEEE Transactions on Software Engineering, SE-6(5):501–502, September/ October 1980. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702767.

See [SW78].

Gordon:1979:MIP

[Gor79a] R. D. Gordon. Measuring improvements in program clarity. IEEE Transactions on Software Engineering, SE-5(2):79– 90, March/April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702601.

Gordon:1979:QJM

[Gor79b] R. D. Gordon. A qualitative justification for a measure of program clarity. IEEE Transactions on Software Engineering, SE-5(2):121–128, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702606.

Ginsburg:1978:DSS

[GR78] S. Ginsburg and E. M. Rounds. Dynamic syntax specification using grammar forms. IEEE Transactions on Software Engineering, SE-4(1):44–55, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702488.

Gries:1976:ICI

[Gri76] D. Gries. An illustration of current ideas on the derivation of correctness proofs and correct programs. IEEE Transactions on Software Engineering, SE-2(4):238–244, December 1976. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702379.

See correction [Gri77].

Gries:1977:CIC

[Gri77] D. Gries. Correction to “An Illustration of Current Ideas on the Derivation of Correctness Proofs and Correct Programs”. IEEE Transactions on Software Engineering, SE-3(3):262, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702435.

See [Gri76].

Gries:1978:MAS

[Gri78]  D. Gries.  The multiple as-

signment statement. IEEE Transactions on Software Engineering, SE-4(2):89–93, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702502.

Gannon:1978:OAF

[GS78] T. F. Gannon and S. D.

Shapiro. An optimal approach to fault tolerant software systems design. IEEE Transactions on Software Engineering, SE-4(5):390–409, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/

                                        stamp.jsp?arnumber=1702554.

See comments [LLS81].

Ghosh:1976:DEM

[GT76] S. P. Ghosh and W. G. Tuel. A design of an experiment to model data base system performance. IEEE Transactions on Software Engineering, SE-2(2):97– 106, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702347.

German:1975:SIA

[GW75] S. M. German and B. Wegbreit. A synthesizer of inductive assertions. IEEE Transactions on Software Engineering, SE-1(1):68–75, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312821.

Gerhart:1976:CSA

[GY76a] S. L. Gerhart and L. Yelowitz. Control structure abstractions of the backtracking programming technique. IEEE Transactions on Software Engineering, SE-2(4):285–292, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702385.

Gerhart:1976:OFA

[GY76b] S. L. Gerhart and L. Yelowitz. Observations of fallibility in applications of modern programming methodologies. IEEE Transactions on Software Engineering, SE-2(3):195–207, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702366.

Hewitt:1979:SPT

[HA79] C. E. Hewitt and R. R. Atkinson. Specification and proof techniques for serializers. IEEE Transactions on Software Engineering, SE-5(1):10–23, January/ February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702583.

Halstead:1979:GES

[Hal79] M. H. Halstead. Guest editorial on software science. IEEE Transactions on Software Engineering, SE-5(2):74–75, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702599.

Hamlet:1977:TPA

[Ham77] R. G. Hamlet. Testing programs with the aid of a compiler. IEEE Transactions on Software Engineering, SE-3(4):279–290, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702444.

Hansen:1977:EMC

[Han77] P. B. Hansen. Experience with modular concurrent programming. IEEE Transactions on Software Engineering, SE-3(2):156–159, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702418.

Hansen:1978:NMP

[Han78] P. B. Hansen. Network: A multiprocessor program. IEEE Transactions on Software Engineering, SE-4(3):194–199, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702521.

Haralick:1977:IAP

[Har77a] R. M. Haralick. Image access protocol for image processing software. IEEE Transactions on Software Engineering, SE-3(2):190–192, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702423.

Harrison:1977:CAV

[Har77b] W. H. Harrison. Compiler analysis of the value ranges for variables. IEEE Transactions on Software Engineering, SE-3(3):243–250, May/ June 1977. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702432.

Harrison:1979:NSC

[Har79]  W. H. Harrison.  A new

strategy for code generation — the general-purpose optimizing compiler. IEEE Transactions on Software Engineering, SE-5(4):367–373, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702642.

Hsiao:1979:DAC

[HKN79] D. K. Hsiao, D. S. Kerr, and Chen-Jen Nee. Database access control in the presence of context dependent protection requirements. IEEE Transactions on Software Engineering, SE-5(4):349–358, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702640.

Howden:1976:RPA

[How76] W. E. Howden. Reliability of the path analysis testing strategy. IEEE Transactions on Software Engineering, SE-2(3):208–215, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702367.

Howden:1977:STD

[How77] W. E. Howden. Symbolic testing and the DISSECT symbolic evaluation system. IEEE Transactions on Software Engineering, SE-3(4):266–278, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702443.

Howden:1978:DSE

[How78a] W. E. Howden. DISSECT — a symbolic evaluation and program testing system. IEEE Transactions on Software Engineering, SE-4(1):70–73, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702491.

Howden:1978:TES

[How78b] W. E. Howden. Theoretical and empirical studies of program testing. IEEE Transactions on Software Engineering, SE-4(4):293–298, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702537.

Hughes:1978:AVE

[HP78] C. E. Hughes and C. P. Pfleeger. ASSIST-V: An environment simulator for IBM 360 systems software development. IEEE Transactions on Software Engineering, SE-4(6):526–530, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702573.

Hewitt:1975:TPA

[HS75] C. E. Hewitt and B. Smith. Towards a programming apprentice. IEEE Transactions on Software Engineering, SE-1(1):26–45, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312818.

Hansen:1978:SIM

[HS78] P. B. Hansen and J. Staunstrup.

Specification and implementation of mutual exclusion. IEEE Transactions on Software Engineering, SE-4(5):365–370, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702551.

Huang:1979:DDF

[Hua79] J. C. Huang. Detection of data flow anomaly through program instrumentation. IEEE Transactions on Software Engineering, SE-5(3):226–236, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702622.

Horning:1977:SHC

[HW77] J. J. Horning and D. B. Wortman. Software hut: A computer program engineering project in the form of a game. IEEE Transactions on Software Engineering, SE-3(4):325–330, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702450.

Hevner:1979:QPD

[HY79] A. R. Hevner and S. B. Yao. Query processing in distributed database system. IEEE Transactions on Software Engineering, SE-5(3):177–187, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702617.

Hamilton:1976:HOS

[HZ76] M. Hamilton and S. Zeldin. Higher order software — a methodology for defining software. IEEE Transactions on Software Engineering, SE-2(1):9–32, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702333.

Irvine:1977:ASE

[IB77] C. A. Irvine and J. W. Brackett. Automated software engineering through structured data management. IEEE Transactions on Software Engineering, SE-3(1):34–40, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702400.

Jones:1976:LEC

[JL76] A. K. Jones and B. H. Liskov.

A language extension for controlling access to shared data. IEEE Transactions on Software Engineering, SE-2(4):277–285, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702384.

Katz:1978:POU

[Kat78] S. Katz. Program optimization using invariants. IEEE Transactions on Software Engineering, SE-4(5):378–389, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702553.

Katzenelson:1979:CDS

[Kat79] J. Katzenelson. Clusters and dialogues for set implementations. IEEE Transactions on Software Engineering, SE-5(3):256– 275, May/June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702625.

Klinger:1977:AP

[Kli77]  A. Klinger.  Artificial pat-

terns. IEEE Transactions on Software Engineering, SE-3(4):301– 306, July/August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702447.

Knuth:1977:DPR

[Knu77] Donald E. Knuth. Deletions that preserve randomness. IEEE Transactions on Software Engineering, SE-3(5):351–359, September/October 1977. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702459.

Kodres:1978:ART

[Kod78a] U. R. Kodres. Analysis of realtime systems by data flowgraphs. IEEE Transactions on Software Engineering, SE-4(3):169– 178, May/June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702518.

Kodres:1978:DSF

[Kod78b] U. R. Kodres. Discrete systems and flowcharts. IEEE Transactions on Software Engineering, SE-4(6):521–525, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702572.

Kieburtz:1978:CM

[KS78] R. B. Kieburtz and A. Silberschatz. Capability managers. IEEE Transactions on Software Engineering, SE-4(6): 467–477, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702566.

Kashyap:1977:AMA

[KSY77] R. L. Kashyap, S. K. C. Subas, and S. B. Yao. Analysis of the multiple-attribute-tree data-base organization. IEEE Transactions on Software Engineering, SE-3(6): 451–467, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702474.

Kundu:1978:NCP

[Kun78] S. Kundu. Note on a constrainedpath problem in program testing. IEEE Transactions on Software Engineering, SE-4(1):75–76, January/February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702493.

Kane:1975:CSF

[KY75] J. R. Kane and S. S. Yau.

Concurrent software fault detection. IEEE Transactions on Software Engineering, SE-1(1):87–99, March 1975. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312823.

Lamport:1977:PCM

[Lam77] L. Lamport. Proving the correctness of multiprocess programs. IEEE Transactions on Software Engineering, SE-3(2):125–143, March/April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.org/
                                        stamp/stamp.jsp?arnumber=1702415.

Lee:1976:APC

[LCC76] R. C. T. Lee, Y. H. Chin, and S. C. Chang. Application of principal component analysis to multikey searching. IEEE Transactions on Software Engineering, SE-2(3):185–193, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702364.

Ledgard:1977:PSN

[Led77] H. F. Ledgard. Production systems: A notation for defining syntax and translation. IEEE Transactions on Software Engineering, SE-3(2):105–124, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702414.

Lin:1979:CPS

[LLD79] W. C. Lin, R. C. T. Lee, and H. C.

Du. Common properties of some multiattribute file systems. IEEE Transactions on Software Engineering, SE-5(2):160–174, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702610.

Lynch:1975:REC

[LLS75] W. C. Lynch, J. W. Langner, and M. S. Schwartz. Reliability experience with Chi/OS. IEEE Transactions on Software Engineering, SE-1(2):253–257, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312845.

Lee:1981:SCC

[LLS81] P. A. Lee, J. L. Lloyd, and S. K.

Shrivastava. Some critical comments on the paper “An Optimal Approach to Fault Tolerant Software Systems Design” by Gannon and Shapiro. IEEE Transactions on Software Engineering, SE-7(6): 608–610, November/December 1981. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702893.

See [GS78].

Lucena:1979:PDU

[LP79] C. J. P. Lucena and T. H. C.

Pequeno. Program derivation using data types: A case study. IEEE Transactions on Software Engineering, SE-5(6): 586–592, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702675.

Liskov:1979:EHC

[LS79] B. H. Liskov and A. Snyder. Exception handling in CLU. IEEE Transactions on Software Engineering, SE-5(6): 546–558, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702672.

Lee:1978:TAA

[LSM78] R. C. T. Lee, J. R. Slagie, and C. T. Mong. Towards automatic auditing of records. IEEE Transactions on Software Engineering, SE-4(5):441–448, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702559.

Liskov:1975:STD

[LZ75] B. H. Liskov and S. N. Zilles. Specification techniques for data abstractions. IEEE Transactions on Software Engineering, SE-1 (1):7–19, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6312816.

McGraw:1979:ACP

[MA79] J. R. McGraw and G. R. Andrews. Access control in parallel programs. IEEE Transactions on Software Engineering, SE-5(1):1–9, January/February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702582.

Marie:1979:AAM

[Mar79] R. A. Marie. An approximate analytical method for general queueing networks. IEEE Transactions on Software Engineering, SE-5(5):530–538, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702663.

Masuda:1979:MMM

[Mas79] T. Masuda. Methods for the measurement of memory utilization and the improvement of program locality. IEEE Transactions on Software Engineering, SE-5(6): 618–631, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702679.

Morgan:1975:CNM

[MBGK75] D. E. Morgan, W. Banks, D. P.

Goodspeed, and R. Kolanko.

A computer network monitoring system. IEEE Transactions on Software Engineering, SE-1(3):299–311, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312855.

Mccammon:1975:ASE

[Mcc75a] S. R. Mccammon. Applied software engineering: A real-time simulator case history. IEEE Transactions on Software Engineering, SE-1(4):377–383, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312868.

Mcclure:1975:TBS

[Mcc75b] C. L. Mcclure. Top-down, bottom-up, and structured programming. IEEE Transactions on Software Engineering, SE-1(4):397–403, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312871.

McCabe:1976:CM

[McC76] T. J. McCabe. A complexity measure. IEEE Transactions on Software Engineering, SE-2(4):308–320, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702388.

Mckeeman:1975:PPL

[Mck75] W. M. Mckeeman. On preventing programming languages from interfering with programming. IEEE Transactions on Software Engineering, SE-1(1):19–26, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312817.

Merwin:1978:GES

[Mer78] R. E. Merwin. Guest editorial: Software management: We must find a way. IEEE Transactions on Software Engineering, SE-4(4):307–308, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702539.

Mills:1976:SD

[Mil76]  H. D. Mills.  Software de-

velopment. IEEE Transactions on Software Engineering, SE-2(4):265–273, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702382.

Misra:1978:AFD

[Mis78a] J. Misra. An approach to formal definitions and proofs of programming principles. IEEE Transactions on Software Engineering, SE-4(5):410–413, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702555.

Misra:1978:SAV

[Mis78b] J. Misra. Some aspects of the verification of loop computations. IEEE Transactions on Software Engineering, SE-4(6): 478–486, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702567.

Misra:1978:TAC

[Mis78c] J. Misra. A technique of algorithm construction on sequences. IEEE Transactions on Software Engineering, SE-4(1):65–69, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702490.

Menasce:1979:LDD

[MM79] D. A. Menasce and R. R. Muntz. Locking and deadlock detection in distributed data bases. IEEE Transactions on Software Engineering, SE-5(3):195–202, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702619.

Moore:1975:IIP

[Moo75] J. S. Moore. Introducing iteration into the Pure Lisp theorem prover. IEEE Transactions on Software Engineering, SE-1(3):328–338, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312857.

Moriconi:1979:DVA

[Mor79] M. S. Moriconi. A designer/ verifier’s assistant. IEEE Transactions on Software Engineering, SE-5(4):387–401, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702644.

Miller:1976:AGF

[MS76] Webb Miller and David L.

Spooner. Automatic generation of floating-point test data. IEEE Transactions on Software Engineering, SE-2(3):223– 226, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702369.

Marcotty:1977:DMS

[MS77] M. Marcotty and F. G. Sayward. The definition mechanism for Standard PL/I. IEEE Transactions on Software Engineering, SE-3(6):416–450, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702473.

Musa:1975:TSR

[Mus75] J. D. Musa. A theory of software reliability and its application. IEEE Transactions on Software Engineering, SE-1(3):312–327, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312856.

Manna:1978:LCP

[MW78a] Z. Manna and R. Waldinger. The logic of computer programming. IEEE Transactions on Software Engineering, SE-4(3):199–229, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702522.

McHenry:1978:SLC

[MW78b] R. C. McHenry and C. E. Walston. Software life cycle management: Weapons process developer. IEEE Transactions on Software Engineering, SE-4(4):334–344, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702543.

Manna:1979:SDR

[MW79] Z. Manna and R. Waldinger.

Synthesis: Dreams → programs. IEEE Transactions on Software Engineering, SE-5(4):294–328, July/August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702636.

Ntafos:1979:PCP

[NH79] S. C. Ntafos and S. L. Hakimi.

On path cover problems in digraphs and applications to program testing. IEEE Transactions on Software Engineering, SE-5(5):520–529, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702662.

Noonan:1975:SPF

[Noo75] R. E. Noonan. Structured programming and formal specification. IEEE Transactions on Software Engineering, SE-1(4):421– 425, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6312875.

Nutt:1977:PPO

[Nut77] G. J. Nutt. A parallel processor operating system comparison. IEEE Transactions on Software Engineering, SE-3(6): 467–475, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702475.

Oldehoeft:1979:DSS

[OB79] R. R. Oldehoeft and L. J. Bass. Dynamic software science with applications. IEEE Transactions on Software Engineering, SE-5(5):497–504, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702659.

Oldehoeft:1977:CBL

[Old77] R. R. Oldehoeft. A contrast between language level measures. IEEE Transactions on Software Engineering, SE-3(6): 476–478, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702476.

Ottenstein:1979:QED

[Ott79]  L. M. Ottenstein.  Quanti-

tative estimates of debugging requirements. IEEE Transactions on Software Engineering, SE-5(5):504–514, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702660.

Paige:1975:PGA

[Pai75] M. R. Paige. Program graphs, an algebra, and their implication for programming. IEEE Transactions on Software Engineering, SE-1(3):286–291, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312853.

Paige:1977:PPG

[Pai77] M. R. Paige. On partitioning program graphs. IEEE Transactions on Software Engineering, SE-3(6): 386–393, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702470.

Parnas:1976:DDP

[Par76] D. L. Parnas. On the design and development of program families. IEEE Transactions on Software Engineering, SE-2(1):1–9, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702332.

Parnas:1979:DSE

[Par79] D. L. Parnas. Designing software for ease of extension and contraction. IEEE Transactions on Software Engineering, SE-5(2):128–138, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702607.

Pederson:1978:KRI

[PB78] J. T. Pederson and J. K. Buckle. Kongsberg’s road to an industrial software methodology. IEEE Transactions on Software Engineering, SE-4(4):263–269, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702533.

Parnas:1976:DSM

[PHW76] D. L. Parnas, G. Handzel, and H. Wurges. Design and specification of the minimal subset of an operating system family. IEEE Transactions on Software Engineering, SE-2(4):301–307, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702387.

Polak:1979:EAP

[Pol79]  W. Polak.  An exercise in

automatic program verification. IEEE Transactions on Software Engineering, SE-5(5):453–458, September/October 1979. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702654.

Pratt:1978:CCD

[Pra78] T. W. Pratt. Control computations and the design of loop control structures. IEEE Transactions on Software Engineering, SE-4(2):81–89, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702501.

Porter:1975:RIP

[PS75] T. Porter and I. Simon. Random insertion into a priority queue structure. IEEE Transactions on Software Engineering, SE-1(3):292–298, September 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312854.

Putnam:1978:GES

[Put78] L. H. Putnam. A general empirical solution to the macro software sizing and estimating problem. IEEE Transactions on Software Engineering, SE-4(4):345–361, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702544.

Randell:1975:SSS

[Ran75] B. Randell. System structure for software fault tolerance. IEEE Transactions on Software Engineering, SE-1(2):220– 232, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312842.

Rubey:1975:QAS

[RDB75] R. J. Rubey, J. A. Dana, and P. W. Biche. Quantitative aspects of software validation. IEEE Transactions on Software Engineering, SE-1(2):150– 155, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312835.

Reisner:1977:UPE

[Rei77]  P. Reisner.  Use of psy-

chological experimentation as an aid to development of a query language. IEEE Transactions on Software Engineering, SE-3(3):218–229, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702430.

Ripley:1978:PSM

[RGH78] G. D. Ripley, R. E. Griswold, and D. R. Hanson. Performance of storage management in an implementation of SNOBOL4. IEEE Transactions on Software Engineering, SE-4(2):130–137, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702506.

Ramamoorthy:1975:TLS

[RH75] C. V. Ramamoorthy and S. B. F. Ho. Testing large software with automated software evaluation systems. IEEE Transactions on Software Engineering, SE-1(1): 46–58, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6312819.

Rose:1976:PCP

[RH76] L. L. Rose and H. Hellerman. Portable character processing in Fortran and fixed integer environments. IEEE Transactions on Software Engineering, SE-2(3):176–185, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702363.

Ramamoorthy:1976:AGP

[RHC76] C. V. Ramamoorthy, S. B. F.

Ho, and W. T. Chen. On the automated generation of program test data. IEEE Transactions on Software Engineering, SE-2(4):293–300, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702386.

Ripley:1977:PPR

[Rip77] G. D. Ripley. Program perspectives: A relational representation of measurement data. IEEE Transactions on Software Engineering, SE-3(4):296–300, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702446.

Ramamoorthy:1975:OPS

[RKC75] C. V. Ramamoorthy, K. H. Kim, and W. T. Chen. Optimal placement of software monitors aiding systematic testing. IEEE Transactions on Software Engineering, SE-1(4):403–411, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312872.

Rypka:1979:DDA

[RL79] D. J. Rypka and A. P. Lucido. Deadlock detection and avoidance for shared logical resources. IEEE Transactions on Software Engineering, SE-5 (5):465–471, September/October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702656.

Robinson:1979:SAT

[Rob79] J. T. Robinson. Some analysis techniques for asynchronous multiprocessor algorithms. IEEE Transactions on Software Engineering, SE-5(1):24–31, January/ February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702584.

Rochkind:1975:SCC

[Roc75] M. J. Rochkind. The source code control system. IEEE Transactions on Software Engineering, SE-1(4):364–370, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312866.

Roman:1977:AFM

[Rom77] G. C. Roman. An argument in favor of mechanized software production. IEEE Transactions on Software Engineering, SE-3(6): 406–415, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702472.

Ross:1977:SAS

[Ros77] D. T. Ross. Structured Analysis (SA): A language for communicating ideas. IEEE Transactions on Software Engineering, SE-3(1):16–34, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702399.

Rosen:1978:TLS

[Ros78] B. K. Rosen. The toy language syndrome. IEEE Transactions on Software Engineering, SE-4(1):76–77, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702494.

Roussopoulos:1979:CCS

[Rou79] N. Roussopoulos. CSDL: A conceptual schema definition language for the design of data base applications. IEEE Transactions on Software Engineering, SE-5(5):481–496, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702658.

Ross:1977:SAR

[RS77] D. T. Ross and K. E. Schoman. Structured analysis for requirements definition. IEEE Transactions on Software Engineering, SE-3(1):6–15, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702398.

Rich:1978:IRL

[RS78] C. Rich and H. E. Shrobe. Initial report on a Lisp programmer’s apprentice. IEEE Transactions on Software Engineering, SE-4(6): 456–467, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702565.

Rowe:1978:ASI

[RT78] L. A. Rowe and F. M. Tonge.

Automating the selection of implementation structures. IEEE Transactions on Software Engineering, SE-4(6):494–506, November/December 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702569.

Ralston:1976:SFE

[RW76] A. Ralston and J. L. Wagener. Structured Fortran — an evolution of Standard Fortran. IEEE Transactions on Software Engineering, SE-2(3):154– 176, September 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702362.

Riddle:1978:BMD

[RWS^{+}78] W. E. Riddle, J. C. Wileden, J. H. Sayler, A. R. Segal, and A. M. Stavely. Behavior modeling during software design. IEEE Transactions on Software Engineering, SE-4(4):283–292, July/ August 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702536.

Reynolds:1976:IBP

[RY76] C. Reynolds and R. T. Yeh. Induction as the basis for program verification. IEEE Transactions on Software Engineering, SE-2(4):244–252, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702380.

Ryder:1979:CCG

[Ryd79] B. G. Ryder. Constructing the call graph of a program. IEEE Transactions on Software Engineering, SE-5(3):216–226, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702621.

Samet:1977:MDF

[Sam77] H. Samet. A machine description facility for compiler testing. IEEE Transactions on Software Engineering, SE-3 (5):343–351, September/October 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702458.

Sholl:1975:SPM

[SB75] H. A. Sholl and T. L. Booth. Software performance modeling using computation structures. IEEE Transactions on Software Engineering, SE-1(4):414–420, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312874.

Shapiro:1977:ESL

[SB77] L. G. Shapiro and R. J. Baron.

ESP^{3}: A language for pattern description and a system for pattern recognition. IEEE Transactions on Software Engineering, SE-3(2):169–183, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702421.

Shrivastava:1978:RRA

[SB78] S. Kumar Shrivastava and J. P.

Banatre. Reliable resource allocation between unreliable processes. IEEE Transactions on Software Engineering, SE-4(3):230–241, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702523.

Schutz:1979:DLP

[Sch79] H. A. Schutz. On the design of a language for programming realtime concurrent processes. IEEE Transactions on Software Engineering, SE-5(3):248–255, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702624.

Soffer:1977:IRB

[SF77] S. B. Soffer and R. B. Friedman. Information retrieval based on Touch-Tone(R) encoded alphanumeric keys. IEEE Transactions on Software Engineering, SE-3(2):184–189, March/ April 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702422.

Schneidewind:1979:CES

[SH79a] N. F. Schneidewind and H. M. Hoffmann. Correction to “An Experiment in Software Error Data Collection and Analysis”. IEEE Transactions on Software Engineering, SE-5(5):538, September/ October 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702664.

See [SH79b].

Schneidewind:1979:ESE

[SH79b] N. F. Schneidewind and H. M. Hoffmann. An experiment in software error data collection and analysis. IEEE Transactions on Software Engineering, SE-5(3):276–286, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702626.

See correction [SH79a].

Shaw:1978:SDF

[Sha78] A. C. Shaw. Software descriptions with flow expressions. IEEE Transactions on Software Engineering, SE-4(3):242–254, May/ June 1978. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702524.

Silberschatz:1979:CSD

[Sil79] A. Silberschatz. Communication and synchronization in distributed systems. IEEE Transactions on Software Engineering, SE-5(6): 542–546, November/December 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702671.

Silberschatz:1977:ECP

[SKB77] A. Silberschatz, R. B. Kieburtz, and A. J. Bernstein. Extending Concurrent Pascal to allow dynamic resource management. IEEE Transactions on Software Engineering, SE-3(3):210– 217, May/June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702429.

Smith:1977:TME

[Smi77] A. J. Smith. Two methods for the efficient analysis of memory address trace data. IEEE Transactions on Software Engineering, SE-3(1):94–101, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702406.

Smith:1978:CSS

[Smi78] A. J. Smith. A comparative study of set associative memory mapping algorithms and their use for cache and main memory. IEEE Transactions on Software Engineering, SE-4(2):121–130, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702505.

Symes:1977:CPS

[SO77] L. R. Symes and R. R. Oldehoeft. Context of problemsolving systems. IEEE Transactions on Software Engineering, SE-3(4):306–309, July/ August 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702448.

Spitzen:1976:SA

[Spi76] J. M. Spitzen. The specification of assemblers. IEEE Transactions on Software Engineering, SE-2(1):33–40, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702334.

Spirn:1979:QNR

[Spi79]  J. R. Spirn.  Queuing net-

works with random selection for service. IEEE Transactions on Software Engineering, SE-5(3):287–289, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702627.

Scott:1975:PPG

[SS75] R. F. Scott and D. B. Simmons. Predicting programming group productivity a communications model. IEEE Transactions on Software Engineering, SE-1(4):411–414, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312873.

Steel:1975:GE

[Ste75] T. B. Steel. Guest editorial. IEEE Transactions on Software Engineering, SE-1(4):349–350, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312863.

Stone:1977:MSA

[Sto77a] H. S. Stone. Multiprocessor scheduling with the aid of network flow algorithms. IEEE Transactions on Software Engineering, SE-3(1):85–93, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702405.

Stoutemyer:1977:CIN

[Sto77b] D. R. Stoutemyer. Computing with impure numbers: Automatic consistency checking and units conversion using computer algebra. IEEE Transactions on Software Engineering, SE-3(6): 478–480, November/December 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702477.

Stone:1978:CLF

[Sto78a] H. S. Stone. Critical load factors in two-processor distributed systems. IEEE Transactions on Software Engineering, SE-4(3):254–258, May/ June 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702525.

Stone:1978:SS

[Sto78b] H. S. Stone. Sorting on STAR. IEEE Transactions on Software Engineering, SE-4(2):138–146, March/April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=1702507.

Stonebraker:1979:CCC

[Sto79] M. Stonebraker. Concurrency control and consistency of multiple copies of data in distributed ingres. IEEE Transactions on Software Engineering, SE-5(3):188–194, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702618.

Stockenberg:1975:SPA

[SV75] J. E. Stockenberg and A. Van Dam. STRUCT programming analysis system. IEEE Transactions on Software Engineering, SE-1(4):384–389, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312869.

Schick:1978:ACS

[SW78] G. J. Schick and R. W. Wolverton. An analysis of competing software reliability models. IEEE Transactions on Software Engineering, SE-4(2):104–120, March/ April 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702504.

See summary [Goe80].

Tai:1978:SEC

[Tai78]  Kuo-Chung Tai.  Syntactic

error correction in programming languages. IEEE Transactions on Software Engineering, SE-4(5):414–425, September/ October 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702556.

Tanenbaum:1976:GPM

[Tan76] A. S. Tanenbaum. A generalpurpose macro processor as a poor man’s compiler-compiler. IEEE Transactions on Software Engineering, SE-2(2):121–125, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702350.

Teichroew:1977:PPC

[TH77] D. Teichroew and E. A. Hershey. PSL/PSA: A computer-aided technique for structured documentation and analysis of information processing systems. IEEE Transactions on Software Engineering, SE-3(1):41–48, January/ February 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702401.

Tompkins:1976:C

[Tom76] H. E. Tompkins. Comments on “Structured Programming in a Production Programming Environment”. IEEE Transactions on Software Engineering, SE-2(1):67, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702338.

See [Bak75].

Towster:1979:CED

[Tow79] E. Towster. A convention for explicit declaration of environments and top-down refinement of data. IEEE Transactions on Software Engineering, SE-5(4):374–386, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702643.

Trivedi:1979:DMC

[TW79] K. S. Trivedi and R. A. Wagner. A decision model for closed queuing networks. IEEE Transactions on Software Engineering, SE-5(4):328–332, July/ August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702637.

VanEmden:1979:PVC

[Van79] M. H. Van Emden. Programming with verification conditions. IEEE Transactions on Software Engineering, SE-5(2):148–159, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702609.

Voigt:1975:PDM

[Voi75] S. Voigt. Program design by a multidisciplinary team. IEEE Transactions on Software Engineering, SE-1(4):370–376, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312867.

Walke:1977:PRB

[Wal77] B. Walke. Program replacement for better throughput. IEEE Transactions on Software Engineering, SE-3(5):369–374, September/October 1977. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1702462.

Waters:1979:MAL

[Wat79] R. C. Waters. A method for analyzing loop programs. IEEE Transactions on Software Engineering, SE-5(3):237–247, May/ June 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=1702623.

Whitten:1975:MCI

[WD75] D. E. Whitten and P. A. D. Demaine. A machine and configuration independent Fortran: Portable Fortran PFortran. IEEE Transactions on Software Engineering, SE-1(1):111–124, March 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

                                        //ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=6312825.

Wegbreit:1975:PEW

[Weg75] B. Wegbreit. Property extraction in well-founded property sets. IEEE Transactions on Software Engineering, SE-1(3):270–285, September 1975. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312852.

Wegbreit:1976:GDP

[Weg76] B. Wegbreit. Goal-directed program transformation. IEEE Transactions on Software Engineering, SE-2(2):69–80, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702344.

Wegbreit:1977:CMP

[Weg77] B. Wegbreit. Constructive methods in program verification. IEEE Transactions on Software Engineering, SE-3(3):193–209, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702428.

Williams:1976:PBS

[WG76] R. Williams and G. M. Giddings. A picture-building system. IEEE Transactions on Software Engineering, SE-2(1):62–66, March 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702337.

Woodward:1979:MCF

[WHH79] M. R. Woodward, M. A. Hennell, and D. Hedley. A measure of control flow complexity in program text. IEEE Transactions on Software Engineering, SE-5(1):45–50, January/ February 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702586.

Wilkes:1976:SES

[Wil76] M. V. Wilkes. Software engineering and structured programming. IEEE Transactions on Software Engineering, SE-2(4):274–276, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702383.

Wirth:1975:APL

[Wir75] N. Wirth. An assessment of the programming language Pascal. IEEE Transactions on Software Engineering, SE-1(2):192– 198, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312839.

Witt:1975:CA

[Wit75] J. Witt. The COLUMBUS approach. IEEE Transactions on Software Engineering, SE-1(4):358–363, December 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312865.

Wulf:1976:ICV

[WLS76] W. A. Wulf, R. L. London, and M. Shaw. An introduction to the construction and verification of Alphard programs. IEEE Transactions on Software Engineering, SE-2(4):253–265, December 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702381.

Woodfield:1979:EUI

[Woo79] S. N. Woodfield. An experiment on unit increase in problem complexity. IEEE Transactions on Software Engineering, SE-5(2):76–79, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702600.

Wortman:1976:SHR

[Wor76] D. B. Wortman. A study of high-resolution timing. IEEE Transactions on Software Engineering, SE-2(2):135–137, June 1976. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702352.

Wortman:1979:LAE

[Wor79] D. B. Wortman. On legality assertions in Euclid. IEEE Transactions on Software Engineering, SE-5(4):359–367, July/ August 1979. CODEN IESEDJ.

ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702641.

Wetherell:1979:TDT

[WS79] C. Wetherell and A. Shannon. Tidy drawings of trees. IEEE Transactions on Software Engineering, SE-5(5):514–520, September/October 1979. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1702661.

Wulf:1975:RHS

[Wul75] W. A. Wulf. Reliable hardware/software architecture. IEEE Transactions on Software Engineering, SE-1(2):233–240, June 1975. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6312843.

Zweben:1979:FDO

[ZH79] S. H. Zweben and M. H. Halstead. The frequency distribution of operators in PL/I programs. IEEE Transactions on Software Engineering, SE-5(2):91–95, March/ April 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702602.

Zelkowitz:1978:ICB

[ZL78] M. V. Zelkowitz and H. J. Larsen. Implementation of a capabilitybased data abstraction. IEEE Transactions on Software Engineering, SE-4(1):56–64, January/ February 1978. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702489.

Zweben:1977:SPS

[Zwe77] S. H. Zweben. A study of the physical structure of algorithms. IEEE Transactions on Software Engineering, SE-3(3):250–258, May/ June 1977. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:

//ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1702433.